### v1

Abstract
Unmanned Aerial Vehicles (UAVs) have become essential tools in inspection and emergency response operations due to their high maneuverability and ability to access hard-to-reach areas. However, their limited battery life significantly restricts their use in long-duration missions. This paper presents a novel tethered marsupial robotic system composed of a UAV and an Unmanned Ground Vehicle (UGV), specifically designed for autonomous, long-duration inspection tasks in Global Navigation Satellite System (GNSS)-denied environments. The system extends the UAV’s operational time by supplying power through a tether connected to high-capacity battery packs carried by the UGV. We detail the hardware architecture based on off-the-shelf components to ensure replicability and describe our full-stack software framework, which is composed of open-source components and built upon the Robot Operating System (ROS). The proposed software architecture enables precise localization using a Direct LiDAR Localization (DLL) method and ensures safe path planning and coordinated trajectory tracking for the integrated UGV–tether–UAV system. We validate the system through three field experiments: (1) a manual flight endurance test to estimate the operational duration, (2) an autonomous navigation test, and (3) an inspection mission to demonstrate autonomous inspection capabilities. Experimental results confirm the robustness and autonomy of the system, its capacity to operate in GNSS-denied environments, and its potential for long-endurance, autonomous inspection and monitoring tasks.
keywords:
GNSS-denied areas, Tethered robots, Long-duration inspection, Marsupial system
1 Introduction
The research and deployment of Unmanned Aerial Vehicles (UAVs) have experienced exponential growth during the twenty-first century in a broad variety of civil applications [1]. Some applications which include UAVs as the main agents are: road traffic monitoring, remote sensing, search and rescue operations in emergency situations, security and surveillance over an area, assessment and spraying in agriculture [2], forest firefighting [3, 4]. In them, UAVs are widely used due to their high maneuverability, mobility, and the decreasing costs of both purchase and maintenance. Other applications that have benefited from the use of UAVs are related to inspection tasks to analyze and monitor civil infrastructure [5] [6]. These tasks use the capabilities of the UAV to move within closed, dangerous, or risky environments, or to reach objectives that are at a great distance. For example, UAVs have been used in structure assessment on buildings and bridges [7], power line inspection, facade and indoor fire monitoring and extinguishing [4, 8] and even sewer inspection [9].
However, a significant limitation of small and medium-sized UAVs is their flight duration, which hinders their ability to conduct comprehensive inspection tasks. In fact, the majority of commercial platforms exhibit flight times of less than thirty minutes, typically measured under laboratory conditions and without any additional payload, such as computing devices or sensors.
There are some alternatives developed for extending the flight duration of UAVs. The most common approach is to perform battery swapping either manually, thus requiring human intervention, or automatically with battery swapping stations [10], requiring the installation of additional infrastructure. Another alternative is to use a swarm of drones [11], which greatly increases the complexity of the system and its maintenance costs. Finally, the use of a tethered UAV fixed to a base station that supplies its power [12] is another approach more related to the one proposed in this paper. In fact, tethered UAVs have gathered the attention of the community [13] with applications as diverse as collaborative load transportation [14], high bandwidth communications, [15, 16] and airborne wind energy [17]. Regrettably, the use of a fixed tether at a given location significantly restricts the operational range of the UAV, thereby preventing it from safely accessing certain areas of interest.
In this paper, we present an autonomous tethered marsupial robotic configuration for long-duration inspection tasks. Our system utilizes an Unmanned Ground Vehicle (UGV) to transport battery packs that supply power through a cable to a UAV equipped with the sensors needed to carry out the inspection. This approach is based on a philosophy of robotic collaboration, offering remarkable advantages over traditional standalone systems by leveraging the unique strengths of each robotic agent [18]. A case of success of a marsupial robotic system can be found in the DARPA 2021 Subterranean Challenge [19], where the CSIRO Data61 Team reached second place [20]. The team used a marsupial system in which a UGV transports a UAV to a take-off point to carry out inspection tasks, which reduces the UAV energy consumption. Another successful case is the Mars 2020 rover and helicopter [21] which is being used for planetary exploration on Mars. In this system, the helicopter augments the capabilities of the rover, exploring large areas faster than the rover, providing reconnaissance on target locations and safe-to-traverse routes. In both applications, the concept of collaboration based on the robotic strengths is used; however, the robots are not tethered to provide energy, and the UAV operation time is still limited by the power it can carry.
An important milestone in the context of marsupial tethered robots is introduced in [22], where the authors describe the development of a tethered UAV system operating alongside a UGV for inspecting stone-mine pillars and assessing structural integrity in underground mines. The UAV is powered via a tether from the UGV, allowing for extended operation without relying solely on its onboard battery. The UGV carries additional batteries, providing continuous power through the tether and effectively addressing the UAV’s flight-time limitations. Despite the robustness of the implemented system, the work focuses on autonomous navigation and inspection around stone pillars, rather than enabling navigation and inspection in general or more complex environments. Furthermore, the article does not provide an analysis of the UAV’s long-term inspection capabilities. A similar collaborative UAV-UGV system connected by a power tether is presented in [12]. This system leverages the strengths of both platforms to extend autonomous navigation in partially mapped environments: the UGV, equipped with a large power supply, powers the UAV, while the UAV contributes aerial surveillance, enhanced situational awareness, and obstacle detection for the UGV. However, this work lacks a detailed description of the system architecture and does not analyze flight duration.
Another approach based on a marsupial system is presented in [23], where the authors introduce an optimal path-planning strategy for a marsupial robotic system consisting of a UGV, a UAV, and a tether connecting both robots. The system follows a sequential navigation strategy through 3D environments, avoiding both ground and aerial obstacles to ensure a collision-free trajectory for the team. Nonetheless, this work does not present a complete analysis of the system’s suitability for autonomous inspection tasks or evaluate its flight endurance. In [24], the authors introduce a collaborative UGV-UAV system for reliable localization and autonomous operation in Global Navigation Satellite System (GNSS)-denied environments. The UGV, equipped with a 3D LiDAR, a fisheye camera, an Ultra-WideBand (UWB) radio, and an Inertal Measurement Unit (IMU), handles mapping and localization, while the UAV uses shared IMU and laser altimeter data from the UGV. An Error-State Extended Kalman Filter fuses these sensor data to estimate the UAV’s relative position, enabling real-time tracking and 3D mapping. However, the system does not consider the tether in its planning process, nor does it address UAV flight-time limitations. In [25], a simulator based on the Robotic Operating System (ROS) version 2 and the Gazebo simulator is presented for marsupial systems where a UAV and a UGV are connected by an adjustable-length hanging tether. This configuration enables realistic modeling of the interactions between the robots and the winch that regulates the tether. Its primary contribution is to offer a tool for validating and testing planning and control algorithms, thereby reducing the costs and risks of physical experiments. However, the work is limited to serving as a tool for testing navigation modules, and it does not facilitate field experiments to quantify the viability of this type of system, the UAV’s flight duration, or even the battery charge status.
This paper extends our previous work presented in [26], in which we presented a short description of the proposed system for long-duration missions. As the extension, this paper gives a detailed description of the main onboard software modules used for localization and navigation purposes. In addition, we present a larger set of validation experiments, including quantitative localization and navigation results. Thus, the main contributions of this paper are as follows:
  1. 1.
    We present our marsupial robotic system’s hardware and software design (see Fig. 1), which considers a UGV tied to a UAV, based on off-the-shelf components in contrast to existing approaches in the literature. In our system, we use a tether that connects the UAV to the UGV. This tether provides the UAV with power, enabling long-term inspection missions.
  2. 2.
    We present the architecture of our software solution. The different software modules, developed in-house, are released as open-source. All the modules have been incrementally validated to reach the final goal of performing autonomous inspection missions with a tethered UAV-UGV marsupial system. This is a significant contribution w.r.t. [26, 27, 28].
  3. 3.
    We pay special emphasis on describing our internally developed localization and navigation systems. In particular, the localization system is based on the one presented in our previous work [27], but is adapted for the current case and for providing multi-robot localization in Section 3. Regarding navigation, we present a reformulation of our trajectory planner from [28], adapted to account for the constraints of the proposed marsupial system presented in this paper.
The contributions of the paper are validated in several field experiments, including an inspection experiment of a civil infrastructure.
The paper is structured as follows: Section 2 provides a basic description of the hardware of the UGV, the UAV and the tether systems, including details on the power system for the UAV. It also presents the basic software architecture of the system. Then, we describe in detail the localization and navigation systems used in autonomous inspection missions in Sections 3 and 4, respectively. Our marsupial system has been extensively validated in experimental tests on three different scenarios. These results are presented in Section 5. Later, we present an extended discussion of our results in Section 6. Finally, the conclusions and future research directions are detailed in Section 7.
Figure 1: Our marsupial system performing autonomous inspection tasks in two buildings with structural damage in the University Pablo de Olavide, Seville (Spain). (a) Detail of the UAV at the abandoned thermal station (Scenario 2). (b) Tethered marsupial system at the old theater building (Scenario 3).
2 System Overview
In this section, we describe the main parts of our marsupial system. The system consists of a UGV, a UAV, and a tether system (see Fig. 2), which provides the UAV with power from additional battery packs onboard the UGV. In this way, we can greatly extend the flight duration of the UAV. The UAV, the UGV and the tether system are described in Sections 2.1, 2.2, and 2.3, respectively. Section 2.4 analyzes the power requirements to calculate the capacity of the batteries on the UGV. Finally, Section 2.5 gives an overview of our software solution for autonomous operation.
Refer to caption Refer to caption
(a) (b)
Figure 2: Marsupial System, UAV tied to UGV. (a) Detail of the UGV equipped with the auxiliary structure that stores the systems of the marsupial configuration. (b) Complete system in an initial position before the UAV takeoff.
2.1 Unmanned Ground Vehicle - UGV
We use an ARCO ground platform from IDMind [29], which is a 4-wheeled holonomic robot with an independent traction system designed for the delivery of heavy loads in factories (see Fig. 3). The platform carries its own batteries, which are two LiFePO4 24V, for motors and electronics. ARCO has an autonomy of 3 hours of operation with a maximum payload of over $100\ kg$ and a maximum speed of 0.8 m/s.
Figure 3: Basic ARCO UGV from IDMind with its onboard batteries.
To carry out the task, we have properly equipped ARCO with the necessary components to achieve the marsupial configuration (see Fig. 2a), which are detailed below:
  • Onboard Computer, model MSI CubiN 8GL-001BEU with an Intel i5 processor.
  • A Sparkfun Razor 9 DoF IMU placed inside of the platform. Used for mapping and localization.
  • LiDAR: A 3D OS1-16 Ouster LiDAR, located on top of the robot. It generates 320k points per second, ranging from 0.2 to 120 meters. This information is used for localization, mapping and obstacle avoidance.
  • Auxiliary structure. Designed primarily to position the tether system at the top of the UGV. It is built from Boch RexRoth aluminum profiles.
  • Battery pack. This is part of the tether power system. Two LiFePO4 12V batteries store the energy that is provided to the UAV to increase its flight time.
  • Inverter. This is part of the tethered power system. It converts the DC energy from the batteries into AC energy for the tether system.
  • Reel System. It is placed on top of the auxiliary structure. It connects the Power System with the UAV.
  • Router. The UGV is equipped with a NIGHTHAWK NetGear AX5400 Router. In our setup, UGV has a wired connection via Ethernet, while the Base Station and the UAV connect via 5G WiFi.
2.2 Unmanned Aerial Vehicle - UAV
Regarding the aerial platform, the model used for the experiments is the DJI Matrice 210 V2 (M210) (see Fig. 4). We selected this platform because of the robustness of the hardware and software, the support that the provider offers, and its ROS integration by means of a Software Development Kit (SDK).
The drone’s structure was adapted in order to carry an i7 NUC computer and a LiDAR, as shown in Fig. 4. The main sensors and systems on the UAV are:
  • M210 onboard sensors. Our UAV includes a single GNSS receiver, an IMU and a barometric altimeter.
  • 3D LiDAR sensor. An OS1-16 Ouster LiDAR is used for environment mapping, localization, and odometry.
  • USB WiFi access point. A D-Link AX1800 Wi-Fi USB Adapter is used to increase the robustness, range and performance of the WiFi connection at the UAV.
Figure 4: M210 UAV and its additional onboard systems.
2.3 Tether System
We use the system LIGH-T V4 Tethered Station (see Fig. 5), provided by the company Elistair [30]. We selected this system due to the support that the company provides and its compatibility with the M210, among other DJI models. The cable of the system ends in an adapter, which is located in the drone battery compartment to provide energy to the drone. Note that a fully charged battery has to be equipped in the UAV in the remaining battery slot on the M210. This battery acts as a backup that is used in case of a disconnection between the UAV and the UGV.
The maximum length of the tether is around seventy meters. Besides, the system works with an AC input. Thus, an inverter is required to connect it to the battery pack. Finally, the reel is connected to a DC motor, which can perform a constant tension to the cable ranging from one to three Newtons.
Figure 5: (a) LIGH-T V4 Tethered Station, including switches, a reel for the cable, and the battery connector. (b) Detail of the battery connector.
2.4 Power System Requirements
We have designed our power system to let the UAV fly up to two hours while with a remaining battery level above 10%. Figure 6 represents the interconnection scheme of the power system, which consists of the following components:
  • LiFePO4 batteries (part of the battery pack) have great characteristics including high energy density, high current discharge, and long durability. Most importantly, they are safer to operate when compared to LiPO batteries [31]. Thus, for powering up the UAV from the UGV, we have selected the Joiry LiFePO4 battery, 12.8 V, 150 Ah, for a total energy of 1920 Wh. Thus, considering the power consumption of the tether system (1200 W), we obtain a theoretical flight duration slightly over one and a half hours. We connected two batteries in parallel to further extend the flight duration (see Fig. 6).
  • We have installed a compact Green Cell® 2000W/4000W inverter from 12V DC to 220V/230V AC. This inverter powers the Elistair system and the router onboard the UGV during the experiment.
  • A backup battery is connected to the UAV in the unused battery slot of the M210. We use the battery model TB55 that provides the UAV with about ten minutes of flight duration in case of failure of the Elistair tether system.
Figure 6: Connection diagram of the power system onboard the UGV and the UAV.
2.5 Software architecture
This section describes the main software components running on each of the platforms and on the laptop used as the Base Station. All of our systems run on ROS Noetic in Ubuntu 20.04. ROS is designed to distribute atomic tasks on each robot in independent processes called ROS nodes. Figure 7 represents the main ROS nodes running on each platform. We detail the tasks carried out by each ROS node below.
All the presented software modules are available in separate repositories at the GitHub profile of the Service Robotics Lab of the University Pablo de Olavide (UPO) [32]. In particular, the marsupial_launchers repository centralizes the high level script files that are used to conduct autonomous marsupial missions. We indicate the repositories of each software module in their description.
2.5.1 Base Station Software
In this section, you can find the main ROS nodes running on the laptop at the Base Station.
  • Trajectory Planner: We use a two-step planner adapted from [28] in which an initial solution computed with an RRT* algorithm is then optimized with a non-linear optimizer that takes into account the kinematic and dynamic constraints and enhances the safety of the routes. This system is described in Sections 4.2 and 4.3. Implementation available in the marsupial_optimizer repository of our GitHub organization [32].
  • Trajectory Coordination: It coordinates the motion of the UAV and UGV (see Section 4.4). controller. Implementation available at the marsupial_mission_interface repository of our GitHub organization [32].
2.5.2 UGV Software
In this section, you can find a list of the most important ROS nodes running at the UGV PC.
  • ROS Master: A ROS Master Core is executed in the UGV computer. It is necessary to establish the communications between the ROS nodes of each computer.
  • ARCO driver: This software interfaces with the hardware board of the robot, and offers standard ROS commands and services for easy integration of the robot with the rest of the system.
  • OS1 LiDAR driver: It communicates the data generated by the sensor with ROS, so that the data is used by other nodes, such as the localization node.
  • UGV Direct LiDAR Localization (DLL): This module uses a prior 3D map of the environment and the LiDAR data to localize the platform using DLL [27]. Please refer to Section 3.2 for details regarding this module. controller. Implementation available at the dll repository of our GitHub organization [32].
  • Trajectory Tracker: This module receives the trajectory computed by the base station and sends commands to ARCO driver module. Details are provided in Section 4.4 for details regarding this system. Implementation available at the arco_traj_tracker repository of our GitHub organization [32].
2.5.3 UAV Software
In this section, you can find a list of the most important ROS nodes running at the UAV PC.
  • DJI driver: Software that interfaces with the hardware board of the UAV, and offers standard ROS topics and services to access the internal DJI devices such as motor and sensors.
  • OS1 LiDAR driver: Same as the UGV.
  • UAV Localization (DLL): Same as the UGV.
  • Trajectory Tracker: Similar to the one at the UGV, but adding a proportional height controller. Implementation available at the matrice_traj_tracker repository of our GitHub organization [32].
Figure 7: Main ROS nodes of the Marsupial System to carry out autonomous missions. The arrows indicate information flow and its main direction. The ROS master node is only needed for establishing the communications between nodes.
3 Localization System
Having a precise localization of the marsupial system is crucial to ensure that the inspection is done as expected during the mission. In outdoor scenarios with good GNSS clearance, we can use Differential or RTK approaches to obtain a precise localization estimation with errors of a few centimeters. However, performing close inspection of large objects (facades, large statues, buildings, etc.) based on GNSS might be subject to significant localization errors (up to several meters) due to the reduction in GNSS satellite visibility and the consequent poor Dilution of Precision (DOP) [33]. LiDAR-based localization approaches have been demonstrated to be reliable and accurate enough to perform long-term robot localization, both aerial and terrestrial.
In this paper, we focus on the localization problem in GNSS-denied areas of our marsupial system. To this end, we use a two-step approach. First, Section 3.1 describes the mapping stage, when we obtain a 3D map of the environment by using a LiDAR sensor. Then, this map is used with the localization method described in Section 3.2 when performing the inspection mission using the same LiDAR sensor.
3.1 Mapping stage
The goal of this stage is to obtain a 3D model of the environment. To this end, a brief manual flight is executed, typically lasting two to five minutes, depending on the spatial extent and complexity of the operational environment. It is important to gather data from all sensors and at different altitudes to ensure proper mapping in the whole scenario. Note that this stage could be omitted as long as a 3D description of the building by means of a CAD file or a Building Information Modeling (BIM) is available, which was not the case in our experimental scenarios.
The purpose of the map is twofold. First, it is used by the localization algorithm as a reference to estimate the pose of both platforms in real time. Second, the 3D map is used to design the inspection plan that should be followed in the experiment. In this paper, we define an inspection plan as a sequence of Points of Interest (PoIs) with orientation that should be visited by the UAV.
We can use any of the state-of-the-art LiDAR-based odometry and SLAM algorithms such as LiDAR Odometry and Mapping (LOAM) [34] or LiDAR-Inertial Odometry and Mapping VE-LIOM [35], to name a few, for experimentally obtaining the 3D map from a sequence of LiDAR and IMU measurements. Our current implementation uses the open source A-LOAM ROS package [36], which is an advanced implementation of the LOAM algorithm that uses Eigen and Ceres Solver to simplify the structure of the software. The mapping results obtained with this package are shown in Fig. 8.
Figure 8: Point clouds of the mapping process using the A-LOAM package. The color represents the height (z-coordinate) of each point. (a) Theater Scenario. (b) Thermal Central Scenario.
3.2 Online Localization
In this section, we assume that we have a 3D point-cloud of the environment, which can be obtained with the methods specified in Section 3.1, and we want to have an online estimation of the pose of our vehicles during the experiment so that the inspection mission is properly executed.
In contrast to [24], our localization system uses a global reference frame shared by both the UAV and UGV, obtaining independent estimations of the pose of both robots with a reprojection error of the order of centimeters (see Section 5). Thus, we follow an approach similar to our previous approach for the MBZIRC 2020 competition [8]. In it, each vehicle used an independent Monte Carlo Localization (MCL) approach integrating a 3D LiDAR sensor, a compass, an altimeter and a GNSS receiver. Even though this system was used with acceptable results, the estimation of the yaw angle of the platform was not very precise, which is crucial when performing inspection tasks.
Hence, we designed an optimization-based alternative, the Direct LiDAR Localization method [27] (DLL). It is a fast direct map-based localization technique using 3D LiDAR for its application. DLL implements a point cloud to map registration based on non-linear optimization of the distance of the points and the map, thus not requiring features, nor point correspondences. Instead, it uses a trilinear interpolation of the Euclidean Distance Field (EDF) representing the environment, which enables enhanced search for the optimal thanks to the continuous gradient computation. The EDF is processed offline during map building; although it can also be generated online with faster implementations like FIESTA [37], significantly reducing computational load compared to kd-tree-based approaches.
This method uses a given initial pose in the 3D map. From it, it is able to estimate the pose of the robot by finding the pose that optimally matches the sensed obstacles in the environment to the prior 3D map. To perform the optimization, it uses the translation estimated by the odometry as the initial guess for the non-linear solver. In particular, we make use of a simple sensor integration for short-term odometry for the UAV in order to save computation time. Regarding the UGV, we use fused odometry from IMU and wheel encoder measurements for odometry. We demonstrate in Section 5 that the DLL algorithm performs much better than 3D MCL methods [38] and achieves comparable or better precision to other map-based localization approaches, such as solutions based on the Iterative Closest Point (ICP) algorithm [39] or the Normal Distributions Transform (NDT) [40], while running one order of magnitude faster.
4 Navigation System
The autonomous navigation system enables the UGV and UAV to perform coordinated navigation in order to safely fulfill an inspection task.
Given the inspection plan as a sequence of PoIs with orientation to be visited, we generate a collision-free trajectory to be tracked for both the UAV and the UGV platforms by means of a two-step trajectory planner described in our previous work [28]. This planner includes a global path planner (Section 4.2) and a non-linear optimizer (Section 4.3) that refines the path into an optimized trajectory. As the Elistair tether system is able to keep the tether taut (see Section 2.3), the planner proposed in [28] has been modified to consider that the length of the cable equals the distance between the UGV and the UAV platform and the tether station. Therefore, only configurations with a direct Line of Sight (LoS) between these platforms are considered feasible. Finally, the computed trajectory is then followed by the trajectory tracker, which is detailed in Section 4.4.
4.1 Trajectory Planning Problem formulation
We define a state in the state space as the combination of the position of the UGV $\mathbf{p}_{g}=(x_{g},y_{g},z_{g})^{T}$, the UAV $\mathbf{p}_{a}=(x_{a},y_{a},z_{a})^{T}$. At any instant, the tether length $l$ is equal to the distance from the UGV to the UAV ($l=\|\mathbf{p}_{a}-\mathbf{p}_{g}\|$). Therefore, in contrast to [28] the length of the tether is not considered as a planner variable but rather considered taut during the whole mission, and thus we have to check for LoS visibility between the platforms to ensure collision-free operation.
The navigation system consists of motion planning which determines the trajectory for the UGV $\mathbf{p}_{g}(t)$, the UAV $\mathbf{p}_{a}(t)$ so that the UAV reaches a given goal position, avoiding obstacles and meeting the constraints of the system. We work on time discrete trajectories (1), forming a set of states for each time step.
$O=\{\mathbf{p}^{i}_{g},\mathbf{p}^{i}_{a},\Delta t^{i}\}_{i=1,...,n}$ (1)
where $i$ is a time step of the trajectory and $n$ is the total number of time steps. $\Delta t^{i}=t^{i}-t^{i-1}$ is the time increment between steps $i$ and $i-1$. This value is the same for UGV and UAV trajectories. For each $\mathbf{p}^{i}_{a}$ , $\mathbf{p}^{i}_{g}$ and $l^{i}$, there is a taut tether configuration $T^{i}$. When needed, we discretize it into a set of $m$ positions $\mathbf{p}_{t}=(x_{t},y_{t},z_{t})$ (2).
$T^{i}=\{\mathbf{p}^{j}_{t}\}_{j=1,...,m}$ (2)
4.2 Path planning
In the first step, the path planner is based on a RRT* algorithm, following the general structure presented in [28] with modifications to accommodate the characteristics of the tethered marsupial system. As we are working with a taut tether, we adapted the algorithms responsible for checking catenary collisions to permit only tether states that ensure LoS between the UGV and UAV platforms, meaning that hanging tether states are not allowed.
We employ the RRT* algorithm [41] to solve the problem due to its ability to deal with high-dimensional spaces. The RRT* plans a path in a six-dimensional space composed of the UGV and UAV positions $\mathbf{x}=\{\mathbf{p}_{g},\mathbf{p}_{a}\}$, from an initial state $\mathbf{x}^{i}$ to a goal state $\mathbf{x}^{g}$ in which only the aerial robot position is set (the final position of the UGV will depend on the UAV goal and the environment). The tether is just evaluated procedurally in the RRT* Steering stage, through the checkTetherFeasibility algorithm, presented in Section 4.2.1. The RRT* cost function is just a weighted sum of the total length of the UAV and UGV paths.
The planner leverages a 3D point cloud to represent the environment for UAV and UGV path planning. Initially, a traversability analysis, similar to the one carried out in [42], identifies traversable areas for the UGV, which are used to generate new positions for the UGV. For the UAV, the full 3D point cloud is organized into a grid containing the EDF, which stores the distances from the grid position to its closest obstacle. We sample the grid points in which the EDF exceeds a safety distance.
Figure 9 shows an example of a path trajectory generated by the path planner. The planner generates a sequence of waypoints for the UAV (represented with a dotted black line) and the UGV (dotted white line) with visibility between the UAV and the UGV at each corresponding waypoint, checked in the checkTetherFeasibility algorithm.
4.2.1 checkTetherFeasibility algorithm
This algorithm is integrated into the RRT* steering process and verifies if a collision-free tether exists between the UGV and UAV. For each new state, the tether length $l^{i}$ is defined as the Euclidean distance between $\mathbf{p}^{i}_{a}$ and $\mathbf{p}^{i}_{g}$. If the tether is feasible, the algorithm marks the corresponding UGV and UAV positions as valid for $l^{i}$, returning true.
Figure 9: Example of path obtained by the proposed path planner based on the RRT* algorithm in the Theater scenario. Points in blue and white color represent explored nodes, i. e. nodes of the RRT*’s tree, for the UAV and the UGV, respectively. Black and white lines represent the obtained path for the UAV and the UGV, respectively. The green point represents the random point generated in the current iteration.
4.3 Trajectory planning
The trajectory planning improves the initial path obtained by the path planner (see previous section) by solving a non-linear optimization problem that considers the time dimension as well as safety and geometrical constraints. The path planning method from the previous section outputs a sequence of collision-free robot positions $\{\mathbf{p}^{i}_{g},\mathbf{p}^{i}_{a}\}_{i=1,...,n}$, where $n$ is the trajectory length. This sequence, however, does not include time-related information as in (1), nor does it consider time or safety constraints.
The method presented in [28] incorporates time information into the initial path to create an initial trajectory solution. To achieve this, the method is spaced equidistantly the waypoints along the path. Then, the $\Delta t^{i}$ values are initialized by assigning constant scalar speeds, $v_{g}$ and $v_{a}$, to the UGV and UAV paths, respectively. Both trajectories use the same $\Delta t^{i}$ to ensure that the UGV and UAV reach each waypoint simultaneously. Thus, the value of $\Delta t^{i}$ for each state in (1) is the largest value between $\|\mathbf{p}^{i}_{g}-\mathbf{p}^{i+1}_{g}\|/v_{g}$ and $\|\mathbf{p}^{i}_{a}-\mathbf{p}^{i+1}_{a}\|/v_{a}$.
As the trajectory is computed through optimization, we retain most constraints from our method proposed in [28]. These include limits on velocity and acceleration, equidistant spacing between consecutive robot poses, maximization of the distance from obstacles for each agent (UGV, UAV, and tether), minimization of trajectory execution time, and maximization of smoothness.
Finally, the optimizer outputs a safer (more distant from obstacles) and smoother trajectory than the initial one, along with speeds and accelerations for traversal, as shown in Fig. 10.
In the particular implementation used in this paper, the primary differences from the base method [28] lie in the tether-related constraints. First, the constraint Unfeasible tether length penalizes tether lengths that exceed the maximum feasible length. Second, for the tether obstacle avoidance constraint, we evaluate only the collision-free LoS between platforms, using a straight-line approximation rather than calculating the catenary. The details of these constraints are presented in Sections 4.3.1 and 4.3.2.
Figure 10: Example of final trajectory computed by our non-linear optimizer. Black line and blue dots represent the path and the trajectory respectively for the UAV. White line and green dots represent the path and the trajectory respectively for the UGV. Red dots represent the tether of the optimized trajectory.
4.3.1 Unfeasible tether length constraint
Equation (5) penalizes the distances $d^{i}_{u}$ between $\mathbf{p}^{i}_{g}$ and $\mathbf{p}^{i}_{a}$ are longer than the maximum feasible length $L_{max}$.
$\delta^{i}_{u}=$$\left\{\begin{array}[]{cc}e^{d^{i}_{u}-L_{max}}-1&,\textrm{if}\ \ d^{i}_{u}\ >\ L_{max}\\ 0&,\textrm{otherwise}\end{array}\right.$ (5)
4.3.2 Tether obstacle avoidance constraint
To check for collisions on the tether, we collect $m$ samples of the tether as expressed in (2). Assuming a taut tether, we have to sample the straight line $l^{i}$ connecting $\mathbf{p}^{i}_{g}$ to $\mathbf{p}^{i}_{a}$. Then, we get the distance to the nearest obstacle of each sample from the EDF, $d^{i}_{ot,j}$. Finally, the residual expressed in (8) is the sum of the inverse of the distances. We increase the weight of those samples closer than a safety distance $\rho_{ot}$ to guarantee higher costs in these cases using $\rho_{j}=\beta$, with $\beta>>1$.
$\delta^{i}_{ot}$ $=$ $\sum_{j=1}^{m}\frac{\rho_{j}}{d^{i}_{ot,j}},\ \rho_{j}=\left\{\begin{array}[]{cc}1&,\mathrm{if}\ \ d^{i}_{ot,j}>\rho_{ot}\\ \beta&,\textrm{otherwise}\end{array}\right.$ (8)
4.4 Trajectory tracking
In this section, we detail the control system designed to track the trajectories generated in Section 4.3. We use a distributed approach in which each platform has its independent trajectory tracker (see Section 4.4.1). Then, a coordination module, detailed in Section 4.4.2, ensures the synchronization of their movements.
4.4.1 UAV and UGV trajectory tracker
We assume that both the UAV and the UGV accept holonomic velocity commands. We have implemented a simple trajectory tracking scheme based on proportional control of each one of the coordinate axes (see Fig. 11). Our trajectory tracker acts similarly in both UAV and UGV platforms, but adding a proportional height controller in the case of the UAV. Finally, we incorporate an additional yaw control in the UAV, which is used when reaching the inspection point to ensure that the vehicle is pointing towards the area of interest.
Figure 11: System diagram of the proposed path tracker.
Regarding the magnitude of the velocity command vector ($\boldsymbol{v}$) in the lateral motion, we make use of trapezoidal velocity profiles for waypoint tracking. The maximum velocity for the trapezoid is set to the planned $\rho_{vg}$ and $\rho_{va}$ respectively, but the actual mean velocity is slightly less than the commanded due to the velocity profile.
Finally, the maximum commanded speed has been limited to match the speed of the slowest subsystem, which in this implementation is the UGV. This has been done for safety reasons and to facilitate the coordination of the vehicles. In the experiments, we set the speeds of both the UGV and the UAV ($\rho_{vg}$ and $\rho_{va}$, respectively) to $0.25m/s$.
4.4.2 UGV and UAV coordination
In order to track the trajectory computed by our approach, the UGV and the UAV should follow the trajectories coordinately. This task is carried out by our Trajectory Coordination Module (TCM) (see Fig. 7). In it, we opt for a loosely-coupled solution based on time synchronization. TCM can only command new waypoints to both the UGV and UAV simultaneously, provided that both platforms have successfully reached the previous one. Thus, when a subsystem reaches its current waypoint, it waits for the other in case of need.
5 Experiments
We carried out different field experiments with our marsupial configuration in three different buildings of the UPO, Seville (Spain). Scenario 1 is located in Building 45, home of the Service Robotic Laboratory facilities. Next, we conducted additional tests in two old, abandoned buildings on the same university campus, both of which show significant structural damage. Scenario 2 is situated in the Abandoned Thermal Station, while Scenario 3 is located in the Old Theater building. Additional details of the experiments are provided in the video attached to this paper.
5.1 Scenario 1: Flight Duration Tests
The main goal of this experiment was to perform the necessary system integrity checks of the whole marsupial system and to estimate the flight endurance of the system in manual mode. To obtain a valid estimation, we mimicked the stop-and-go behavior in an inspection mission, obtaining a total trajectory length exceeding one kilometer on both the UAV and the UGV. Figure 12 shows two snapshots obtained during the experiment.
Refer to caption Refer to caption
Figure 12: Two snapshots of the flight duration tests at building 45.
The experiment exceeded one hour duration, which allowed us to get information about the discharging rate of both the onboard UAV and the batteries on the UGV used for powering the UAV via the Elistair tether system. Table 1 represents the level of the batteries powering the tether system and the onboard UAV battery used as a backup, for ease of comparison, during the experiment. In the second row, we can find the battery percentage of the onboard battery used for backup purposes on the UAV. Note that this battery slowly discharges until stabilizing at 86%, when the voltage level of the battery reaches the one provided by the Elistair system. Regarding the batteries onboard the UGV, their discharge rate remains constant during the experiment and very close together. At the end of the experiment, the remaining battery level was $53\%$, giving an average power consumption on the UAV of 1635 W.
Table 1: Charge of the batteries, in [%] of the different systems as the mission time progresses.
Battery Percentage
Time [min] 0 15 30 45 60
Onboard UAV Battery [%] 100 92 88 86 86
Battery Pack 1 [%] 100 88 76 64 52
Battery Pack 2 [%] 100 89 77 66 54
5.2 Scenario 2: UAV Autonomous Navigation Test
In Scenario 2, we visited the Abandoned Thermal Station building of the UPO. This building was built in the decade of the 1950s and presented great structural damage and debris on its surface that prevented our UGV ARCO platform from safely performing autonomous navigation on it. Therefore, we designed some experiments focused on the UAV system to test the indoor UAV localization and navigation systems.
5.2.1 UAV Localization tests
In a first set of experiments, we tested the ability of our localization system to provide us with an online estimation of the pose of the UAV in manual mode. To this end, we designed a challenging experiment in which we made our UAV move around the whole experimental scenario at different height levels while recording the data from the sensors. Then, we tested different localization methods to compare their results and their average computation time. To evaluate the precision of the localization algorithm, we got the distance from each LiDAR detection inside a measurement to the closest obstacle in the map, to check for errors in the localization. The lower the average error in a measurement is, the better the vehicle is localized in the environment. This was necessary due to the lack of ground truth from external systems available. Table 2 represents the mean distance from the LiDAR points to the closest obstacle with different localization methods and their average execution time. Results indicate that our DLL method used in our system outperforms the MCL in terms of precision and runs significantly faster when compared to the ICP method, achieving similar errors. Moreover, the measured distance from the closest obstacle results indicate that the localization error should be in the order of centimeters, while also inferring a good precision in the estimation of the yaw. This precision is enough to safely perform cooperative missions in our marsupial system. Thus, DLL is a very convenient alternative for having precise localization when performing long-endurance inspection missions.
Table 2: Average computation time of the update step on each method and mean distance from each point of the point-cloud (PC) to the closest obstacle in the map in Scenario 2.
Method Avg. Computation time (s) Avg. PC distance to obstacle (m)
AMCL-3D 0.54 0.45
ICP 2.67 0.13
DLL 0.08 0.14
5.2.2 UAV autonomous flight tests
Once we demonstrated that our localization system was able to obtain a precise estimation of the pose of the robot in real-time, we executed several autonomous tests for the UAV to test our trajectory tracking system. To this end, a simple inspection plan was designed and generated the trajectory to be tracked by the UAV by using our trajectory planner, assuming a static UGV. Figure 13 represents the generated trajectories to perform the UAV autonomous navigation test represented in the 3D model of the environment. Our system was able to autonomously follow the trajectory, obtaining mean lateral cross-track and height errors of 0.07 m and 0.08 m, respectively.
Figure 13: Reference trajectory (blue line) tracking experiments in Scenario 2. The position of the reel system is on a green dot and the red lines represent the tether configurations during the experiment.
5.3 Scenario 3: Emulated Inspection Test
Finally, to test the whole system in autonomous mode, we emulated a structural assessment inspection. In Scenario 3, our marsupial system is configured with a mission designed to look for emulated defects in the Old Theater building of the UPO. To emulate the defects, we disposed of an array of twelve AR markers with a side of fifteen centimeters on one of the walls of the building (see Fig. 1b). We equipped the UAV with a high-resolution camera to automatically detect the markers by using the ArUco library [43].
Figure 8.a represents an obtained 3D map of the environment, which was used for localization and path planning purposes on both platforms in Scenario 3. The goal of the experiment is to make an in-depth inspection of one of the walls of the building. To this end, we disposed of the PoIs of our inspection plan following a zig-zag line. Then, we generated a trajectory for the marsupial system using our trajectory planner module (see Section 4.3).
In the experimental session, the batteries of the platform allowed us to perform the execution three times without the need of changing any batteries, for a total flight time of about fifty minutes. The trajectories commanded to the UAV and the UGV are represented in Fig. 14. The UAV platform accurately tracked the trajectories, detecting all emulated defects in three experimental runs, as illustrated in Fig. 15. Statistics on the error in the estimated position of the detected ArUco markers are provided in Table 3. In particular, the mean detection error of each marker is on the order of centimeters. Moreover, the cases with greater errors are usually related to a low number of observations of the AR marker.
Figure 14: Computed trajectory for the inspection mission of UGV (in green dots) and UAV (in blue dots). The square and the sphere in white color represent the initial and the final positions, respectively.
Figure 15: Current and past detections of emulated defects obtained in the inspection mission of Scenario 3.
Table 3: Average error and number of observations in the position estimation of the AR markers.
AR ID’s 411 412 413 415 416 417 418 420 421 422 423 477
Error (cm) 4.2 12.6 13.6 19.8 9.8 6.2 8.7 11.6 18.7 3.1 13.0 7.6
Observations 88 28 277 28 215 445 198 568 344 307 205 29
6 Discussion and take-home messages
This paper presents the design, implementation, and validation of a marsupial robotic system capable of performing long-duration inspections in GNSS-denied environments. Experimental results confirm the hypothesis that a tethered UAV-UGV system can significantly extend UAV flight time, opening the door to long-duration inspections. The accuracy in detecting AR markers, with a mean error in the order of centimeters, demonstrates the viability of the system for detailed inspection tasks.
Even though power consumption was higher than expected, the platform proved capable of performing inspection missions lasting more than two hours. The additional discharge may be due to extra connected components which were not included in the calculations, such as the LiDAR and the router, or due to heat losses in the different voltage conversions. However, the flight duration can be further extended by adding more battery packs to the UGV, as we are far away from its payload limits.
A summary of the most valuable lessons we have learned from experimentation is mentioned in the Section 6.1. Then, Section 6.2 lists the advantages and disadvantages of our approach compared to other alternatives.
6.1 Lessons Learned
Several factors are essential for the success of a multi-robot experiment. In this section, we highlight some key insights that we have acquired from our experimentation.
  • Design a setup procedure and a checklist to ensure experimental quality. When performing experiments with complex systems such as our tether marsupial system, it is important to carefully design a procedure for initializing and checking the systems before starting a mission. We follow the steps outlined in Algorithm 1. In Steps 1–3, the UGV is started system first, as its onboard computer hosts the ROS master node required for system-wide communication. Additionally, we power up the UAV as late as possible (Step 8) to maximize its available flight time. Conducting integrity tests on each onboard sensor is essential to avoid incomplete experiment recordings. Specifically, we verify that each sensor’s data rate and quality meet expected standards—this is done in Step 4 for the UGV and Step 10 for the UAV.
  • Avoid dangerous configurations of the UAV, the UGV and the tether. The Elistair LIGH-T V4 station has a DC motor that performs a constant tension to the tether, keeping it taut. Experimentally, we found out that a tension of one Newton is enough to keep it straight in most situations. The main challenge here is that this tension introduces disturbances that must be managed by the UAV’s control system. When tension is applied downwards, it has a similar effect to increasing the system’s payload. However, when tension is applied laterally, the control system must adjust the UAV’s attitude to compensate for the disturbances. Therefore, it is advisable to penalize configurations with low elevation angles. Lastly, we secure the tether as close to the center of gravity of the UAV as possible to prevent the generation of torques.
  • Prepare and label several UAV batteries the day before an experimental session. UAVs need a significant amount of energy to complete long-duration flights, which typically requires fresh batteries for each experimental run. However, LiPO batteries should not be stored fully charged for extended periods. Therefore, it is advisable to check and charge the UAV batteries the day before an experimental session. It also applies to our system because even if the energy consumption of the onboard UAV backup battery is low during the experiment (see Table 1), the M210 firmware expects the onboard batteries to be fully charged to power up the system.
Algorithm 1 Experiment Setup and Execution Checklist
1:Turn on the battery packs and the inverter at the UGV.
2:Turn on router onboard the UGV.
3:Turn on the UGV and start its ROS nodes, including the ROS master process necessary for communications.
4:Turn on laptop at the base station and check for ROS connectivity with the UGV.
5:Perform integrity tests of the sensors and software systems on the UGV.
6:Set the UAV and the UGV at their initial poses.
7:Turn on the Elistair Power Tether.
8:Just before the experiment: Turn on the UAV computer.
9:Check for connectivity to the UAV system
10:Perform integrity tests to the sensors and software systems onboard the UAV.
11:Start the TCM module on the base station.
12:Start the recording of the experiments including video and data logging.
13:Start the mission.
6.2 Marsupial systems: benefits and drawbacks
With all the knowledge gathered from experimentation, we are confident that our marsupial system is a viable alternative that has multiple advantages over existing approaches in the literature. Thus, the following list summarizes the benefits of our approach.
  • The UGV can supply power to the UAV via the cable, significantly extending flight time compared to the UAV’s limited battery capacity. UAV time fly can extend over one hour.
  • Unlike UAVs tethered to a fixed location, the UGV provides mobility, enabling the UAV to inspect a wider range of areas of interest.
  • The UAV offers aerial mobility and monitoring capabilities, while the UGV provides a stable power source and additional payload capacity.
  • The cable can serve as a reliable communication link and can also enhance safety by providing a physical connection between the two robots.
  • Marsupial UAV-UGV systems have proven valuable in hazardous and confined environments, as exemplified by their use by the CSIRO team during the DARPA Subterranean Challenge. Our system has been able to carry out a long-endurance inspection task, in a GNSS-denied environment, with accurate localization and safe trajectory planning.
However, even though our approach has been demonstrated as an effective alternative to perform long-duration inspection tasks, there are some disadvantages compared to other existing approaches. Besides, our approach has room for improvement in some areas.
  • The tether can introduce significant disturbances in the UAV control subsystem. Interactions between the tether and the UAV generate dynamic forces that affect stability and control.
  • The tether restricts the UAV’s range of motion and requires careful management to prevent collisions. Maintaining proper tension is crucial: downward tension impacts the system payload, while lateral tension requires UAV control adjustments to compensate for perturbations.
  • There is a risk of the tether becoming entangled or snagged on obstacles.Therefore, the mission should be carefully planned to prevent the cable from touching any obstacles. In addition, the cable should also monitored during the experiments, which should be canceled in the event of entanglements or other unexpected events.
  • Variable speed trajectory tracking can be implemented to smoothen the coordinated movement of the UAV and the UGV.
7 Conclusions
This paper presents our solution for long-duration inspection: a tethered marsupial system connecting a UAV and a UGV. In it, we detail the hardware and software configurations that allow autonomous navigation in UAV inspection tasks. Notably, the system achieved over one hour of flight time—and up to a maximum of two hours—thanks to the power supplied by the batteries onboard the UGV. Therefore, the proposed marsupial system successfully extends the operational capabilities of the UAV.
The experimental results obtained across three different scenarios validate the robustness and operational capabilities of our system in real-world conditions, delivering reliable inspection performance in challenging, GNSS-denied environments. Specifically, a series of tests conducted at the University Pablo de Olavide confirmed the system’s ability to carry out autonomous, long-duration inspection missions. Notably, our implementation of the DLL localization system provided real-time pose estimations with centimeter-level accuracy, outperforming state-of-the-art methods. Furthermore, our path tracking system successfully followed the reference trajectories, maintaining maximum tracking errors below ten centimeters. As a result, the system was able to detect simulated defects with centimeter-level accuracy during the emulated inspection test.
Future work will focus on refining the power system to further extend the operational range. Additionally, we are developing a continuous control system within the TCM module to enable smooth coordination between platforms by adjusting the tracking velocity of each system, moving away from the current Stop & Go approach. Moreover, we are interested in expanding the inspection capabilities of the system by integrating additional sensors required for in-depth inspection tasks, such as radar and IR cameras. However, the UAV’s payload limitations may prevent it from carrying all sensors simultaneously. To address this, we are taking two key actions. First, we are designing modular sensor attachments that allow easy reconfiguration of the UAV based on the specific requirements of each inspection mission. Second, Elistair offers a tethered station which includes an optical fiber communication link. In this way, we could remove the PC onboard the UAV, making room for installing additional sensors.
Declarations
[…]
Funding This work was supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21, funded by MCIN/AEI/ 10.13039/501100011033 and ”European Union NextGenerationEU/PRTR”.
[…]
Conflict of interest/Competing interests The authors declare that they have no conflicts of interest or competing interests.
[…]
Ethics Approval Not Applicable.
[…]
Consent for publication Not Applicable.
[…]
Consent to participate Not Applicable.
[…]
Data availability
Data will be available upon reasonable request. Also, the code is available on the GitHub organization "robotics-upo" [32].
[…]
Authors’ Contributions Simón Martínez-Rozas and David Alejo: Conceptualization, data curation, software, investigation, methodology, analysis, experiments, writing original draft. José Javier Carpio: Mechanical development and maintenance of the robot, experiments, writing-review and editing. Fernando Caballero and Luis Merino: Conceptualization, validation, resources, supervision, funding acquisition, project management, writing-review and editing.
[…]
Acknowledgements
We would like to thank the staff of the Service Robotics Laboratory at the Universidad Pablo de Olavide for their unselfish help during the experiments. We would also like to give special thanks to Sara Victoria Hernández Díaz for suggesting different experimental scenarios for structural inspection and granting us access to them.

### v2

@@
 However, their limited battery life significantly restricts their use in long-duration missions.
-This paper presents a novel tethered marsupial robotic system composed of a UAV and an Unmanned Ground Vehicle (UGV), specifically designed for autonomous, long-duration inspection tasks in Global Navigation Satellite System (GNSS)-denied environments.
+This paper presents a tethered marsupial robotic system composed of a UAV and an Unmanned Ground Vehicle (UGV), specifically designed for autonomous, long-duration inspection tasks in Global Navigation Satellite System (GNSS)-denied environments.
 The system extends the UAV’s operational time by supplying power through a tether connected to high-capacity battery packs carried by the UGV.
-We detail the hardware architecture based on off-the-shelf components to ensure replicability and describe our full-stack software framework, which is composed of open-source components and built upon the Robot Operating System (ROS).
+Our work details the hardware architecture based on off-the-shelf components to ensure replicability and describes our full-stack software framework used by the system, which is composed of open-source components and built upon the Robot Operating System (ROS).
 The proposed software architecture enables precise localization using a Direct LiDAR Localization (DLL) method and ensures safe path planning and coordinated trajectory tracking for the integrated UGV–tether–UAV system.
-We validate the system through three field experiments: (1) a manual flight endurance test to estimate the operational duration, (2) an autonomous navigation test, and (3) an inspection mission to demonstrate autonomous inspection capabilities.
-Experimental results confirm the robustness and autonomy of the system, its capacity to operate in GNSS-denied environments, and its potential for long-endurance, autonomous inspection and monitoring tasks.
-keywords:
-GNSS-denied areas, Tethered robots, Long-duration inspection, Marsupial system
+We validate the system through three sets of field experiments involving (i) three manual flight endurance tests to estimate the operational duration, (ii) three experiments for validating the localization and the trajectory tracking systems, and (iii) three executions of an inspection mission to demonstrate autonomous inspection capabilities.
+The results of the experiments confirm the robustness and autonomy of the system in GNSS-denied environments.
+Finally, all experimental data have been made publicly available to support reproducibility and to serve as a common open dataset for benchmarking.
 1 Introduction
-The research and deployment of Unmanned Aerial Vehicles (UAVs) have experienced exponential growth during the twenty-first century in a broad variety of civil applications [1].
-Some applications which include UAVs as the main agents are: road traffic monitoring, remote sensing, search and rescue operations in emergency situations, security and surveillance over an area, assessment and spraying in agriculture [2], forest firefighting [3, 4].
+Research and deployment of Unmanned Aerial Vehicles (UAVs) have experienced exponential growth during the 21st century in a broad variety of civil applications [1].
+Some applications that include UAVs as the main agents are road traffic monitoring, remote sensing, search and rescue operations in emergency situations, security and surveillance over an area, assessment and spraying in agriculture [2], forest firefighting [3].
 In them, UAVs are widely used due to their high maneuverability, mobility, and the decreasing costs of both purchase and maintenance.
-Other applications that have benefited from the use of UAVs are related to inspection tasks to analyze and monitor civil infrastructure [5] [6].
-These tasks use the capabilities of the UAV to move within closed, dangerous, or risky environments, or to reach objectives that are at a great distance.
-For example, UAVs have been used in structure assessment on buildings and bridges [7], power line inspection, facade and indoor fire monitoring and extinguishing [4, 8] and even sewer inspection [9].
+Other applications that have benefited from the use of UAVs are related to inspection tasks to analyze and monitor civil infrastructure [4, 5].
+These tasks use the capabilities of the UAV to move within closed, dangerous, or risky environments or to reach objectives that are at a great distance.
+For example, UAVs have been used in the evaluation of the structure of buildings and bridges [6], the inspection of power lines, the monitoring and extinguishing of fires in the facade and interior of buildings [3].
 However, a significant limitation of small and medium-sized UAVs is their flight duration, which hinders their ability to conduct comprehensive inspection tasks.
-In fact, the majority of commercial platforms exhibit flight times of less than thirty minutes, typically measured under laboratory conditions and without any additional payload, such as computing devices or sensors.
-There are some alternatives developed for extending the flight duration of UAVs.
+In fact, the majority of commercial platforms exhibit flight times of less than thirty minutes, typically measured under laboratory conditions and without any additional payload, such as computing devices or sensors [7].
+Fully equipped autonomous drones carrying the necessary sensor payload seldom achieve flight times of more than twenty minutes [8].
+There are some alternatives to develop long-duration inspection tasks.
 The most common approach is to perform battery swapping either manually, thus requiring human intervention, or automatically with battery swapping stations [9], requiring the installation of additional infrastructure.
-Another alternative is to use a swarm of drones [11], which greatly increases the complexity of the system and its maintenance costs.
-Finally, the use of a tethered UAV fixed to a base station that supplies its power [12] is another approach more related to the one proposed in this paper.
-In fact, tethered UAVs have gathered the attention of the community [13] with applications as diverse as collaborative load transportation [14], high bandwidth communications, [15, 16] and airborne wind energy [17].
-Regrettably, the use of a fixed tether at a given location significantly restricts the operational range of the UAV, thereby preventing it from safely accessing certain areas of interest.
+Another alternative is to use a team of robots flying either alternately or simultaneously, so that the inspection mission can be carried out without interruptions or in a shorter amount of time [10].
+However, the use of a team of UAVs greatly increases the complexity of the system and its acquisition and maintenance costs.
+Finally, tethered UAVs connected to a base station that supplies them with power have attracted the attention of the research community [11, 12] with applications as diverse as collaborative load transportation [13], high-bandwidth communications [14, 15] and airborne wind energy [16].
+Regrettably, the use of a fixed tether at a specific location significantly limits the operational range of the UAV, thereby preventing safe access to certain areas of interest.
 In this paper, we present an autonomous tethered marsupial robotic configuration for long-duration inspection tasks.
@@
 This approach is based on a philosophy of robotic collaboration, offering remarkable advantages over traditional standalone systems by leveraging the unique strengths of each robotic agent [17].
-A case of success of a marsupial robotic system can be found in the DARPA 2021 Subterranean Challenge [19], where the CSIRO Data61 Team reached second place [20].
-The team used a marsupial system in which a UGV transports a UAV to a take-off point to carry out inspection tasks, which reduces the UAV energy consumption.
-Another successful case is the Mars 2020 rover and helicopter [21] which is being used for planetary exploration on Mars.
-In this system, the helicopter augments the capabilities of the rover, exploring large areas faster than the rover, providing reconnaissance on target locations and safe-to-traverse routes.
-In both applications, the concept of collaboration based on the robotic strengths is used; however, the robots are not tethered to provide energy, and the UAV operation time is still limited by the power it can carry.
-An important milestone in the context of marsupial tethered robots is introduced in [22], where the authors describe the development of a tethered UAV system operating alongside a UGV for inspecting stone-mine pillars and assessing structural integrity in underground mines.
+A case of success of a marsupial robotic system can be found in the Defense Advanced Research Projects Agency (DARPA) 2021 Subterranean Challenge [18], where the CSIRO Data61 Team reached second place [19].
+The team used a marsupial system in which a UGV transports a UAV to a take-off point to carry out inspection tasks, which reduces the UAV’s energy consumption.
+Another successful case is the Mars 2020 rover and helicopter [20] which are being used for planetary exploration on Mars.
+In this system, the helicopter augments the capabilities of the rover, exploring large areas faster than the rover and providing reconnaissance on target locations and safe-to-traverse routes.
+In both applications, the concept of collaboration based on the robotic strengths is used; however, the robots are not tethered to provide energy, so the UAV operation time is still limited by the power it can carry.
+Following the tethered marsupial approach, an important milestone is introduced in [21], where the authors describe the development of a tethered UAV system operating alongside a UGV for inspecting stone-mine pillars and assessing structural integrity in underground mines.
 The UAV is powered via a tether from the UGV, allowing for extended operation without relying solely on its onboard battery.
@@
 In [24], the authors introduce a collaborative UGV-UAV system for reliable localization and autonomous operation in Global Navigation Satellite System (GNSS)-denied environments.
-The UGV, equipped with a 3D LiDAR, a fisheye camera, an Ultra-WideBand (UWB) radio, and an Inertal Measurement Unit (IMU), handles mapping and localization, while the UAV uses shared IMU and laser altimeter data from the UGV.
-An Error-State Extended Kalman Filter fuses these sensor data to estimate the UAV’s relative position, enabling real-time tracking and 3D mapping.
+The UGV, equipped with a 3D LiDAR, a fisheye camera, an Ultra-WideBand (UWB) radio, and an Inertial Measurement Unit (IMU), handles mapping and localization, while the UAV uses shared IMU and laser altimeter data from the UGV.
+An Error-State Extended Kalman Filter fuses this sensor data to estimate the UAV’s relative position, enabling real-time tracking and 3D mapping.
 However, the system does not consider the tether in its planning process, nor does it address UAV flight-time limitations.
 In [25], a simulator based on the Robotic Operating System (ROS) version 2 and the Gazebo simulator is presented for marsupial systems where a UAV and a UGV are connected by an adjustable-length hanging tether.
-This configuration enables realistic modeling of the interactions between the robots and the winch that regulates the tether.
+This configuration enables realistic modeling of the interactions between the robots and the winch that regulates the length of the tether.
 Its primary contribution is to offer a tool for validating and testing planning and control algorithms, thereby reducing the costs and risks of physical experiments.
@@
 This paper extends our previous work presented in [26], in which we presented a short description of the proposed system for long-duration missions.
-As the extension, this paper gives a detailed description of the main onboard software modules used for localization and navigation purposes.
-In addition, we present a larger set of validation experiments, including quantitative localization and navigation results.
+As an extension, this paper gives a detailed description of the main onboard software modules used for control, localization and navigation purposes.
+Moreover, we present here extended field reports that show that our proposed system can perform autonomous inspection tasks in cluttered environments, which are still a challenge in tethered UAV-UGV systems [12].
+To this end, we present a set of validation experiments, including quantitative localization and navigation results.
 Thus, the main contributions of this paper are as follows:
   1. 1.
-    We present our marsupial robotic system’s hardware and software design (see Fig.
-1), which considers a UGV tied to a UAV, based on off-the-shelf components in contrast to existing approaches in the literature.
-In our system, we use a tether that connects the UAV to the UGV.
-This tether provides the UAV with power, enabling long-term inspection missions.
+    We present the hardware and software design of the proposed marsupial robotic system as a whole (see Figure 1).
+It consists of a UGV linked to a UAV and is built using off-the-shelf components in contrast to existing approaches in the literature [21].
+In our system, we use a cable that connects the UAV to the UGV.
+This tether supplies the UAV with power, enabling long-term inspection missions.
+To the best of our knowledge, this is the first system to demonstrate a continuous operation exceeding two hours.
   2. 2.
@@
 All the modules have been incrementally validated to reach the final goal of performing autonomous inspection missions with a tethered UAV-UGV marsupial system.
-This is a significant contribution w.r.t.
-[26, 27, 28].
+This is a significant contribution with respect to [26, 27, 28].
   3. 3.
-    We pay special emphasis on describing our internally developed localization and navigation systems.
+    We pay special emphasis on describing our developed localization and navigation systems.
 In particular, the localization system is based on the one presented in our previous work [27], but is adapted for the current case and for providing multi-robot localization in Section 3.
-Regarding navigation, we present a reformulation of our trajectory planner from [28], adapted to account for the constraints of the proposed marsupial system presented in this paper.
-The contributions of the paper are validated in several field experiments, including an inspection experiment of a civil infrastructure.
+Regarding navigation, we present a reformulation of our trajectory planner from [28], adapted to account for the constraints of the proposed marsupial system.
+In particular, in this paper we always consider a straight tether instead of the variable-length hanging tether approach presented in [28].
 The paper is structured as follows: Section 2 provides a basic description of the hardware of the UGV, the UAV and the tether systems, including details on the power system for the UAV.
@@
 Then, we describe in detail the localization and navigation systems used in autonomous inspection missions in Sections 3 and 4, respectively.
-Our marsupial system has been extensively validated in experimental tests on three different scenarios.
+Our marsupial system has been extensively validated in experimental tests in three different scenarios.
 These results are presented in Section 5.
@@
 Finally, the conclusions and future research directions are detailed in Section 7.
-Figure 1: Our marsupial system performing autonomous inspection tasks in two buildings with structural damage in the University Pablo de Olavide, Seville (Spain).
+Refer to caption Refer to caption
+(a) (b)
+Figure 1: Our marsupial system performing autonomous inspection tasks in two buildings with structural damage in the University Pablo de Olavide (UPO), Seville (Spain).
 (a) Detail of the UAV at the abandoned thermal station (Scenario 2).
 (b) Tethered marsupial system at the old theater building (Scenario 3).
-2 System Overview
-In this section, we describe the main parts of our marsupial system.
-The system consists of a UGV, a UAV, and a tether system (see Fig.
-2), which provides the UAV with power from additional battery packs onboard the UGV.
-In this way, we can greatly extend the flight duration of the UAV.
-The UAV, the UGV and the tether system are described in Sections 2.1, 2.2, and 2.3, respectively.
-Section 2.4 analyzes the power requirements to calculate the capacity of the batteries on the UGV.
+2 Materials and Methods
+In this section, we describe the main components of our marsupial system.
+The system consists of a UGV, a UAV, and a tether system (see Figure 2), which supplies the UAV with power from additional battery packs onboard the UGV.
+This configuration allows us to significantly extend the flight duration of the UAV.
+The UGV, the UAV and the tether system are described in Sections 2.1, 2.2 and 2.3, respectively.
+Section 2.4 analyzes the power requirements used to determine the capacity of the batteries on the UGV.
 Finally, Section 2.5 gives an overview of our software solution for autonomous operation.
@@
 (a) Detail of the UGV equipped with the auxiliary structure that stores the systems of the marsupial configuration.
-(b) Complete system in an initial position before the UAV takeoff.
-2.1 Unmanned Ground Vehicle - UGV
-We use an ARCO ground platform from IDMind [29], which is a 4-wheeled holonomic robot with an independent traction system designed for the delivery of heavy loads in factories (see Fig.
-3).
-The platform carries its own batteries, which are two LiFePO4 24V, for motors and electronics.
-ARCO has an autonomy of 3 hours of operation with a maximum payload of over $100\ kg$ and a maximum speed of 0.8 m/s.
+(b) Marsupial system in an initial position before the experiment.
+2.1 Unmanned Ground Vehicle (UGV)
+We use an ARCO ground platform from IDMind, which is a four-wheeled holonomic robot with an independent traction system designed for the delivery of heavy loads in factories (see Figure 3).
+The base platform has two batteries for motors and electronics, respectively, that give ARCO an autonomy of three hours of operation with a maximum payload of over one hundred kilograms and a maximum speed of 0.8 m/s.
 Figure 3: Basic ARCO UGV from IDMind with its onboard batteries.
-To carry out the task, we have properly equipped ARCO with the necessary components to achieve the marsupial configuration (see Fig. 2a), which are detailed below:
-  • Onboard Computer, model MSI CubiN 8GL-001BEU with an Intel i5 processor.
-  • A Sparkfun Razor 9 DoF IMU placed inside of the platform. Used for mapping and localization.
-  • LiDAR: A 3D OS1-16 Ouster LiDAR, located on top of the robot.
-It generates 320k points per second, ranging from 0.2 to 120 meters.
-This information is used for localization, mapping and obstacle avoidance.
-  • Auxiliary structure. Designed primarily to position the tether system at the top of the UGV. It is built from Boch RexRoth aluminum profiles.
-  • Battery pack. This is part of the tether power system. Two LiFePO4 12V batteries store the energy that is provided to the UAV to increase its flight time.
-  • Inverter. This is part of the tethered power system. It converts the DC energy from the batteries into AC energy for the tether system.
-  • Reel System. It is placed on top of the auxiliary structure. It connects the Power System with the UAV.
-  • Router. The UGV is equipped with a NIGHTHAWK NetGear AX5400 Router. In our setup, UGV has a wired connection via Ethernet, while the Base Station and the UAV connect via 5G WiFi.
-2.2 Unmanned Aerial Vehicle - UAV
-Regarding the aerial platform, the model used for the experiments is the DJI Matrice 210 V2 (M210) (see Fig.
-4).
+Additionally, we have properly equipped ARCO with the necessary components to achieve the marsupial configuration (see Figure 2a).
+Most importantly, the ARCO platform includes an internal computer that runs the internal software responsible for processing data from its IMU and 3D LiDAR sensors to localize the platform in real time.
+It also provides the rest of the system with high-speed 5G Wi-Fi connectivity via its integrated router.
+Finally, the tether system and additional batteries for powering the UAV are installed on the platform.
+All components of the ARCO platform are listed in Table 1.
+Table 1: Main components of the ARCO UGV platform.
+Component Model Quantity
+Onboard Computer MSI Cubin 8GL-001BEU (i5-11400) 1
+IMU Sparkfun Razor 9 DoF 1
+LiDAR Ouster OS1-16 1
+UGV Internal Battery LiFePO 75 Ah@24 V 2
+UAV Power Battery Joiry LiFePO 150 Ah@12 V 2
+Inverter Green Cell 2000 W/4000 W 1
+Tether System Elistair LIGH-T V4 Tethered Station (LTS4) 1
+WiFi Router NIGHTHAWK NetGear AX5400 1
+2.2 Unmanned Aerial Vehicle (UAV)
+Regarding the aerial platform, the model used for the experiments is the DJI Matrice 210 V2 (M210) (see Figure 4).
 We selected this platform because of the robustness of the hardware and software, the support that the provider offers, and its ROS integration by means of a Software Development Kit (SDK).
-The drone’s structure was adapted in order to carry an i7 NUC computer and a LiDAR, as shown in Fig. 4. The main sensors and systems on the UAV are:
-  • M210 onboard sensors. Our UAV includes a single GNSS receiver, an IMU and a barometric altimeter.
-  • 3D LiDAR sensor. An OS1-16 Ouster LiDAR is used for environment mapping, localization, and odometry.
-  • USB WiFi access point. A D-Link AX1800 Wi-Fi USB Adapter is used to increase the robustness, range and performance of the WiFi connection at the UAV.
+The main additions are its onboard computer and a LiDAR sensor, as shown in Figure 4.
+In addition, we have installed an external access point to increase the throughput and robustness of the WiFi connectivity (see Table 2).
 Figure 4: M210 UAV and its additional onboard systems.
+Table 2: Main components added to the UAV platform
+Component Model
+Onboard computer NUC8i7BEH 32GB RAM
+LiDAR Ouster OS1-16
+Internal battery DJI TB55 (backup)
+WiFi Access Point D-Link AX1800 Wi-Fi USB Adapter
 2.3 Tether System
-We use the system LIGH-T V4 Tethered Station (see Fig.
-5), provided by the company Elistair [30].
-We selected this system due to the support that the company provides and its compatibility with the M210, among other DJI models.
-The cable of the system ends in an adapter, which is located in the drone battery compartment to provide energy to the drone.
+We use a LIGH-T V4 Tethered Station (LTS4), provided by the company Elistair [29] (see Figure 5) to power the UAV, due to its compatibility with the M210, among other DJI models.
+The cable of the system ends in an adapter, which is located in one of the two drone battery slots to provide energy to the drone.
+The maximum length of the cable is around seventy meters.
 Note that a fully charged battery has to be equipped in the UAV in the remaining battery slot on the M210.
-This battery acts as a backup that is used in case of a disconnection between the UAV and the UGV.
-The maximum length of the tether is around seventy meters.
-Besides, the system works with an AC input.
-Thus, an inverter is required to connect it to the battery pack.
-Finally, the reel is connected to a DC motor, which can perform a constant tension to the cable ranging from one to three Newtons.
-Figure 5: (a) LIGH-T V4 Tethered Station, including switches, a reel for the cable, and the battery connector. (b) Detail of the battery connector.
+This battery acts as a backup that is used in case of a disconnection between the UAV and the LTS4.
+The cable is connected to a DC motor, which can perform a constant tension on the cable ranging from one to three Newtons.
+Figure 5: (a) LIGH-T V4 Tethered Station (LTS4), including switches, a reel for the cable, and the battery connector. (b) Detail of the battery connector.
 2.4 Power System Requirements
-We have designed our power system to let the UAV fly up to two hours while with a remaining battery level above 10%.
+We have designed our power system to allow the UAV to fly up to two hours with a remaining battery level above 10%.
 Figure 6 represents the interconnection scheme of the power system, which consists of the following components:
-  • LiFePO4 batteries (part of the battery pack) have great characteristics including high energy density, high current discharge, and long durability.
+  • We have installed a compact Green Cell® 2000 W/4000 W inverter from 12 V DC to 220 V/230 V AC, with an effectiveness of 85%.
+This inverter powers the LTS4 and the internal UGV router during the experiment.
+  • We use the Joiry LiFePO4 batteries, which have a capacity of 150 Ah at a nominal voltage of 12.8 V.
+With one battery we obtain a theoretical flight duration slightly over seventy minutes, taking into account the effectiveness of the inverter.
+Thus, we connect two batteries in parallel to further extend the flight duration (see Figure 6).
+LiFePO4 batteries have great characteristics, including high energy density, high current discharge, and long durability.
 Most importantly, they are safer to operate when compared to LiPO batteries [30].
-Thus, for powering up the UAV from the UGV, we have selected the Joiry LiFePO4 battery, 12.8 V, 150 Ah, for a total energy of 1920 Wh.
-Thus, considering the power consumption of the tether system (1200 W), we obtain a theoretical flight duration slightly over one and a half hours.
-We connected two batteries in parallel to further extend the flight duration (see Fig.
-6).
-  • We have installed a compact Green Cell® 2000W/4000W inverter from 12V DC to 220V/230V AC. This inverter powers the Elistair system and the router onboard the UGV during the experiment.
-  • A backup battery is connected to the UAV in the unused battery slot of the M210.
-We use the battery model TB55 that provides the UAV with about ten minutes of flight duration in case of failure of the Elistair tether system.
+  • A backup battery is connected to the UAV in the remaining battery slot of the M210.
+We use the battery model TB55 that provides the UAV with about ten minutes of flight duration in case of failure of the LTS4.
+This time is sufficient for the monitoring operator to take control of the UAV and land it safely in the event of an LTS4 system malfunction.
 Figure 6: Connection diagram of the power system onboard the UGV and the UAV.
 2.5 Software Architecture
-This section describes the main software components running on each of the platforms and on the laptop used as the Base Station.
-All of our systems run on ROS Noetic in Ubuntu 20.04.
-ROS is designed to distribute atomic tasks on each robot in independent processes called ROS nodes.
-Figure 7 represents the main ROS nodes running on each platform.
-We detail the tasks carried out by each ROS node below.
-All the presented software modules are available in separate repositories at the GitHub profile of the Service Robotics Lab of the University Pablo de Olavide (UPO) [32].
-In particular, the marsupial_launchers repository centralizes the high level script files that are used to conduct autonomous marsupial missions.
-We indicate the repositories of each software module in their description.
-2.5.1 Base Station Software
-In this section, you can find the main ROS nodes running on the laptop at the Base Station.
-  • Trajectory Planner: We use a two-step planner adapted from [28] in which an initial solution computed with an RRT* algorithm is then optimized with a non-linear optimizer that takes into account the kinematic and dynamic constraints and enhances the safety of the routes.
-This system is described in Sections 4.2 and 4.3.
-Implementation available in the marsupial_optimizer repository of our GitHub organization [32].
-  • Trajectory Coordination: It coordinates the motion of the UAV and UGV (see Section 4.4).
-controller.
-Implementation available at the marsupial_mission_interface repository of our GitHub organization [32].
-2.5.2 UGV Software
-In this section, you can find a list of the most important ROS nodes running at the UGV PC.
-  • ROS Master: A ROS Master Core is executed in the UGV computer. It is necessary to establish the communications between the ROS nodes of each computer.
-  • ARCO driver: This software interfaces with the hardware board of the robot, and offers standard ROS commands and services for easy integration of the robot with the rest of the system.
-  • OS1 LiDAR driver: It communicates the data generated by the sensor with ROS, so that the data is used by other nodes, such as the localization node.
-  • UGV Direct LiDAR Localization (DLL): This module uses a prior 3D map of the environment and the LiDAR data to localize the platform using DLL [27].
-Please refer to Section 3.2 for details regarding this module.
-controller.
-Implementation available at the dll repository of our GitHub organization [32].
-  • Trajectory Tracker: This module receives the trajectory computed by the base station and sends commands to ARCO driver module.
-Details are provided in Section 4.4 for details regarding this system.
-Implementation available at the arco_traj_tracker repository of our GitHub organization [32].
-2.5.3 UAV Software
-In this section, you can find a list of the most important ROS nodes running at the UAV PC.
-  • DJI driver: Software that interfaces with the hardware board of the UAV, and offers standard ROS topics and services to access the internal DJI devices such as motor and sensors.
-  • OS1 LiDAR driver: Same as the UGV.
-  • UAV Localization (DLL): Same as the UGV.
-  • Trajectory Tracker: Similar to the one at the UGV, but adding a proportional height controller. Implementation available at the matrice_traj_tracker repository of our GitHub organization [32].
-Figure 7: Main ROS nodes of the Marsupial System to carry out autonomous missions.
-The arrows indicate information flow and its main direction.
-The ROS master node is only needed for establishing the communications between nodes.
+The proposed marsupial system relies on a distributed software architecture implemented entirely in ROS Noetic running on Ubuntu 20.04.
+ROS provides the natural framework for decomposing the system into independent modules, each implemented as a ROS node, which can run on the UGV, the UAV, or the laptop used as the base station.
+Figure 7 illustrates the overall architecture, showing the nodes allocated to each platform.
+The functionality of these nodes is described in detail in the following sections.
+All software modules are openly available through the GitHub profile of the Service Robotics Lab, Universidad Pablo de Olavide (UPO) ( https://github.com/robotics-upo, accessed on 3 Nov 2025), ensuring reproducibility and transparency.
+Figure 7: Main ROS nodes of the marsupial system used to execute autonomous missions are distributed across the UAV computer (in blue), the Base Station (in green), and the UGV (in red).
+The arrows indicate the flow of information and its predominant direction.
+The ROS master node is required only to establish communication between the nodes.
+At the base station, two key modules orchestrate the system.
+The Trajectory Planner employs a two-step procedure adapted from [28]: first, an initial solution is generated with a Rapidly exploring Random Tree Star (RRT*), and then refined using a non-linear optimizer that incorporates the kinematic and dynamic constraints of both robots while improving safety margins.
+The second module, the Trajectory Coordinator Module (TCM), ensures the synchronized motion of the UAV and UGV during joint missions.
+Onboard the UGV, the architecture begins with the ROS Master, which establishes the communication backbone for all distributed nodes.
+The robot’s hardware is accessed through the ARCO driver, which exposes standard ROS commands and services.
+Environmental perception is provided by the 3D LiDAR driver, whose data feeds directly into the Direct LiDAR Localization (DLL) module [27], responsible for localization against a prior 3D map (see Section 3.2 for more details).
+Trajectory execution is managed by the Trajectory Tracker, which translates planned paths into actionable commands for the ARCO driver (see Section 4.4 for more details).
+The UAV runs a similar software stack, with specific adaptations for aerial operation.
+The DJI driver interfaces with the UAV’s onboard hardware, exposing sensors and actuators through ROS topics and services.
+Like the UGV, the UAV also runs the 3D LiDAR driver and the DLL module, ensuring accurate positioning even in GNSS-denied environments.
+Finally, the UAV executes trajectories using its own trajectory tracker, similar in functionality to the UGV’s tracker but extended with a proportional controller for altitude regulation.
+Altogether, this modular and distributed software architecture enables robust cooperation between the UAV and UGV, leveraging open-source ROS modules to provide a reproducible framework for marsupial robotics research.
 3 Localization System
-Having a precise localization of the marsupial system is crucial to ensure that the inspection is done as expected during the mission.
-In outdoor scenarios with good GNSS clearance, we can use Differential or RTK approaches to obtain a precise localization estimation with errors of a few centimeters.
-However, performing close inspection of large objects (facades, large statues, buildings, etc.) based on GNSS might be subject to significant localization errors (up to several meters) due to the reduction in GNSS satellite visibility and the consequent poor Dilution of Precision (DOP) [33].
+Having a precise localization of the marsupial system is crucial to ensure that the inspection is performed as expected during the mission.
+In outdoor scenarios with good GNSS coverage, we can use Differential or Real-Time Kinematic (RTK) approaches to obtain a precise localization estimation with errors of a few centimeters.
+However, performing close inspections of large objects (facades, large statues, buildings, etc.) based on GNSS might be subject to significant localization errors (up to several meters) due to the reduction in GNSS satellite visibility and the consequent poor Dilution of Precision (DOP) [31].
 LiDAR-based localization approaches have been demonstrated to be reliable and accurate enough to perform long-term robot localization, both aerial and terrestrial.
@@
 In this paper, we define an inspection plan as a sequence of Points of Interest (PoIs) with orientation that should be visited by the UAV.
-We can use any of the state-of-the-art LiDAR-based odometry and SLAM algorithms such as LiDAR Odometry and Mapping (LOAM) [34] or LiDAR-Inertial Odometry and Mapping VE-LIOM [35], to name a few, for experimentally obtaining the 3D map from a sequence of LiDAR and IMU measurements.
-Our current implementation uses the open source A-LOAM ROS package [36], which is an advanced implementation of the LOAM algorithm that uses Eigen and Ceres Solver to simplify the structure of the software.
-The mapping results obtained with this package are shown in Fig.
-8.
+We can use any of the state-of-the-art LiDAR-based odometry and Simultaneous Localization and Mapping (SLAM) algorithms, such as LiDAR Odometry and Mapping (LOAM) [32] or Versatile and Efficient LiDAR-Inertial Odometry and Mapping (VE-LIOM) [33], to name a few, for experimentally obtaining the 3D map from a sequence of LiDAR and IMU measurements.
+Our current implementation uses the open-source advanced implementation LOAM (A-LOAM) ROS package [34], which is an advanced implementation of the LOAM algorithm that uses Eigen and Ceres Solver to simplify the structure of the software.
+The mapping results obtained with this package are shown in Figure 8.
 Figure 8: Point clouds of the mapping process using the A-LOAM package. The color represents the height (z-coordinate) of each point. (a) Theater scenario. (b) Thermal central scenario.
@@
 In this section, we assume that we have a 3D point-cloud of the environment, which can be obtained with the methods specified in Section 3.1, and we want to have an online estimation of the pose of our vehicles during the experiment so that the inspection mission is properly executed.
-In contrast to [24], our localization system uses a global reference frame shared by both the UAV and UGV, obtaining independent estimations of the pose of both robots with a reprojection error of the order of centimeters (see Section 5).
-Thus, we follow an approach similar to our previous approach for the MBZIRC 2020 competition [8].
-In it, each vehicle used an independent Monte Carlo Localization (MCL) approach integrating a 3D LiDAR sensor, a compass, an altimeter and a GNSS receiver.
-Even though this system was used with acceptable results, the estimation of the yaw angle of the platform was not very precise, which is crucial when performing inspection tasks.
-Hence, we designed an optimization-based alternative, the Direct LiDAR Localization method [27] (DLL).
-It is a fast direct map-based localization technique using 3D LiDAR for its application.
-DLL implements a point cloud to map registration based on non-linear optimization of the distance of the points and the map, thus not requiring features, nor point correspondences.
+We have designed an optimization-based localization system, DLL [27].
+It is a fast, direct, map-based localization technique using 3D LiDAR for its application.
+DLL implements a point cloud to map registration based on non-linear optimization of the distance of the points and the map, thus not requiring features nor point correspondences.
 Instead, it uses a trilinear interpolation of the Euclidean Distance Field (EDF) representing the environment, which enables enhanced search for the optimal thanks to the continuous gradient computation.
-The EDF is processed offline during map building; although it can also be generated online with faster implementations like FIESTA [37], significantly reducing computational load compared to kd-tree-based approaches.
+The EDF is processed offline during map building, although it can also be generated online with faster implementations like FIESTA [35], significantly reducing computational load compared to kd-tree-based approaches.
 This method uses a given initial pose in the 3D map.
@@
 Regarding the UGV, we use fused odometry from IMU and wheel encoder measurements for odometry.
-We demonstrate in Section 5 that the DLL algorithm performs much better than 3D MCL methods [38] and achieves comparable or better precision to other map-based localization approaches, such as solutions based on the Iterative Closest Point (ICP) algorithm [39] or the Normal Distributions Transform (NDT) [40], while running one order of magnitude faster.
+We demonstrate in Section 5 that the DLL algorithm performs much better than the 3D Monte Carlo Localization (MCL) method [36, 37] and achieves comparable or better precision to other map-based localization approaches, such as solutions based on the Iterative Closest Point (ICP) algorithm [38] or the Normal Distributions Transform (NDT) [39], while running one order of magnitude faster.
 4 Navigation System
-The autonomous navigation system enables the UGV and UAV to perform coordinated navigation in order to safely fulfill an inspection task.
-Given the inspection plan as a sequence of PoIs with orientation to be visited, we generate a collision-free trajectory to be tracked for both the UAV and the UGV platforms by means of a two-step trajectory planner described in our previous work [28].
-This planner includes a global path planner (Section 4.2) and a non-linear optimizer (Section 4.3) that refines the path into an optimized trajectory.
-As the Elistair tether system is able to keep the tether taut (see Section 2.3), the planner proposed in [28] has been modified to consider that the length of the cable equals the distance between the UGV and the UAV platform and the tether station.
-Therefore, only configurations with a direct Line of Sight (LoS) between these platforms are considered feasible.
-Finally, the computed trajectory is then followed by the trajectory tracker, which is detailed in Section 4.4.
+The autonomous navigation system enables the UGV and UAV to perform coordinated navigation in order to safely execute an inspection task.
+Given the inspection plan as a sequence of PoIs with orientation to be visited, we generate a collision-free trajectory to be tracked for both the UAV and the UGV platforms by using a modification of the planner described in our previous work [28].
+Note that this approach generates trajectories in which all the components of the system (UAV, UGV and tether) do not collide with any obstacles.
+While some approaches may allow the cable to collide with obstacles [40], we have experimentally found that these collisions may lead to unexpected entanglements of the cable with small obstacles (e.g., screws, hooks, etc.).
+These entanglements are difficult to detect, and planning and executing the required unraveling maneuvers is even more challenging.
+Therefore, we avoid any contact between the tether and obstacles to ensure safe autonomous navigation of our marsupial system.
+The proposed planner includes a global path planner (Section 4.2) and a non-linear optimizer (Section 4.3) that refines the path into an optimized trajectory.
+As the LTS4 system is able to keep the tether taut (see Section 2.3), the planner proposed in [28] has been modified to consider that at any moment the length of the cable equals the distance between the UGV and the UAV platform and LTS4.
+Therefore, only configurations with a direct Line of Sight (LoS) between these platforms are considered collision-free.
+Finally, the computed trajectory is followed by the trajectory tracker, which is detailed in Section 4.4.
 4.1 Trajectory Planning Problem Formulation
@@
 Therefore, in contrast to [28] the length of the tether is not considered as a planner variable but rather considered taut during the whole mission, and thus we have to check for LoS visibility between the platforms to ensure collision-free operation.
-The navigation system consists of motion planning which determines the trajectory for the UGV $\mathbf{p}_{g}(t)$, the UAV $\mathbf{p}_{a}(t)$ so that the UAV reaches a given goal position, avoiding obstacles and meeting the constraints of the system.
-We work on time discrete trajectories (1), forming a set of states for each time step.
+Our motion planning algorithm determines the trajectory for the UGV $\mathbf{p}_{g}(t)$, the UAV $\mathbf{p}_{a}(t)$ so that the UAV reaches a given goal position, avoiding obstacles.
+We work on time-discrete trajectories (1), forming a set of states for each time step $i$.
 $O=\{\mathbf{p}^{i}_{g},\mathbf{p}^{i}_{a},\Delta t^{i}\}_{i=1,...,n}$ (1)
 where $i$ is a time step of the trajectory and $n$ is the total number of time steps.
-$\Delta t^{i}=t^{i}-t^{i-1}$ is the time increment between steps $i$ and $i-1$.
+$\Delta t^{i}=t^{i}-t^{i-1}$ is the time increment between steps $i$ and i1i-1.
 This value is the same for UGV and UAV trajectories.
@@
 4.2 Path Planning
-In the first step, the path planner is based on a RRT* algorithm, following the general structure presented in [28] with modifications to accommodate the characteristics of the tethered marsupial system.
+In the first step, the path planner is based on an RRT* algorithm, following the general structure presented in [28] with modifications to accommodate the characteristics of the tethered marsupial system.
 As we are working with a taut tether, we adapted the algorithms responsible for checking catenary collisions to permit only tether states that ensure LoS between the UGV and UAV platforms, meaning that hanging tether states are not allowed.
 We employ the RRT* algorithm [41] to solve the problem due to its ability to deal with high-dimensional spaces.
-The RRT* plans a path in a six-dimensional space composed of the UGV and UAV positions $\mathbf{x}=\{\mathbf{p}_{g},\mathbf{p}_{a}\}$, from an initial state $\mathbf{x}^{i}$ to a goal state $\mathbf{x}^{g}$ in which only the aerial robot position is set (the final position of the UGV will depend on the UAV goal and the environment).
+The RRT* generates a path in a six-dimensional space composed of the UGV and UAV positions $\mathbf{x}=\{\mathbf{p}_{g},\mathbf{p}_{a}\}$, from an initial state $\mathbf{x}^{i}$ to a goal state $\mathbf{x}^{g}$ in which only the aerial robot position is set (the final position of the UGV will depend on the UAV goal and the environment).
 The tether is just evaluated procedurally in the RRT* Steering stage, through the checkTetherFeasibility algorithm, presented in Section 4.2.1.
@@
 Initially, a traversability analysis, similar to the one carried out in [42], identifies traversable areas for the UGV, which are used to generate new positions for the UGV.
-For the UAV, the full 3D point cloud is organized into a grid containing the EDF, which stores the distances from the grid position to its closest obstacle.
-We sample the grid points in which the EDF exceeds a safety distance.
+For the UAV, we compute the EDF from the full 3D point cloud, obtaining a grid representation that stores the distances from the grid position to its closest obstacle.
+New samples of the position of the UAV are obtained randomly on the grid points in which the EDF exceeds a safety distance.
 Figure 9 shows an example of a path trajectory generated by the path planner.
-The planner generates a sequence of waypoints for the UAV (represented with a dotted black line) and the UGV (dotted white line) with visibility between the UAV and the UGV at each corresponding waypoint, checked in the checkTetherFeasibility algorithm.
+The planner generates a sequence of waypoints for the UAV (represented with a dotted black line) and the UGV (dotted white line) with direct visibility between the UAV and the UGV at each corresponding waypoint, checked in the checkTetherFeasibility algorithm.
+Figure 9: Example of the path obtained by the proposed path planner based on the RRT* algorithm in the Theater scenario.
+Points in blue and white color represent explored nodes, i.e., nodes of the RRT*’s tree, for the UAV and the UGV, respectively.
+Black and white lines represent the obtained path for the UAV and the UGV, respectively.
+The green point represents the random point generated in the current iteration.
 4.2.1 checkTetherFeasibility Algorithm
@@
 If the tether is feasible, the algorithm marks the corresponding UGV and UAV positions as valid for $l^{i}$, returning true.
-Figure 9: Example of path obtained by the proposed path planner based on the RRT* algorithm in the Theater scenario.
-Points in blue and white color represent explored nodes, i.
-e.
-nodes of the RRT*’s tree, for the UAV and the UGV, respectively.
-Black and white lines represent the obtained path for the UAV and the UGV, respectively.
-The green point represents the random point generated in the current iteration.
 4.3 Trajectory Planning
@@
 The method presented in [28] incorporates time information into the initial path to create an initial trajectory solution.
-To achieve this, the method is spaced equidistantly the waypoints along the path.
+To achieve this, the method spaces the waypoints equidistantly along the path.
 Then, the $\Delta t^{i}$ values are initialized by assigning constant scalar speeds, $v_{g}$ and $v_{a}$, to the UGV and UAV paths, respectively.
@@
 These include limits on velocity and acceleration, equidistant spacing between consecutive robot poses, maximization of the distance from obstacles for each agent (UGV, UAV, and tether), minimization of trajectory execution time, and maximization of smoothness.
-Finally, the optimizer outputs a safer (more distant from obstacles) and smoother trajectory than the initial one, along with speeds and accelerations for traversal, as shown in Fig. 10.
+Finally, the optimizer outputs a safer (more distant from obstacles) and smoother trajectory than the initial one, along with speeds and accelerations for traversal, as shown in Figure 10.
 In the particular implementation used in this paper, the primary differences from the base method [28] lie in the tether-related constraints.
-First, the constraint Unfeasible tether length penalizes tether lengths that exceed the maximum feasible length.
+First, the Unfeasible Tether Length Constraint penalizes tether lengths that exceed the maximum feasible length.
 Second, for the Tether Obstacle Avoidance Constraint, we evaluate only the collision-free LoS between platforms, using a straight-line approximation rather than calculating the catenary.
-The details of these constraints are presented in Sections 4.3.1 and 4.3.2.
+The details of these constraints are presented in Sections 4.3.1 and 4.3.2, respectively.
 Figure 10: Example of final trajectory computed by our non-linear optimizer.
-Black line and blue dots represent the path and the trajectory respectively for the UAV.
-White line and green dots represent the path and the trajectory respectively for the UGV.
+The black line and blue dots represent the path and the trajectory, respectively, for the UAV.
+The white line and green dots represent the path and the trajectory, respectively, for the UGV.
 Red dots represent the tether of the optimized trajectory.
@@
 Assuming a taut tether, we have to sample the straight line $l^{i}$ connecting $\mathbf{p}^{i}_{g}$ to $\mathbf{p}^{i}_{a}$.
-Then, we get the distance to the nearest obstacle of each sample from the EDF, $d^{i}_{ot,j}$.
+Then, we obtain the distance to the nearest obstacle of each sample from the EDF, $d^{i}_{ot,j}$.
 Finally, the residual expressed in (8) is the sum of the inverse of the distances.
@@
 In this section, we detail the control system designed to track the trajectories generated in Section 4.3.
-We use a distributed approach in which each platform has its independent trajectory tracker (see Section 4.4.1).
-Then, a coordination module, detailed in Section 4.4.2, ensures the synchronization of their movements.
-4.4.1 UAV and UGV trajectory tracker
-We assume that both the UAV and the UGV accept holonomic velocity commands.
-We have implemented a simple trajectory tracking scheme based on proportional control of each one of the coordinate axes (see Fig.
-11).
-Our trajectory tracker acts similarly in both UAV and UGV platforms, but adding a proportional height controller in the case of the UAV.
+We use a distributed approach in which each platform has its independent trajectory tracker.
+Then, a coordination module ensures the synchronization of their movements.
+Thus, for trajectory tracking, we assume that both the UAV and the UGV accept holonomic velocity commands.
+We have implemented a simple trajectory tracking scheme based on proportional control of each one of the coordinate axes (see Figure 11).
+Our trajectory tracker acts similarly in both UAV and UGV platforms but adds a proportional height controller in the case of the UAV.
 Finally, we incorporate an additional yaw control in the UAV, which is used when reaching the inspection point to ensure that the vehicle is pointing towards the area of interest.
@@
 Regarding the magnitude of the velocity command vector ($\boldsymbol{v}$) in the lateral motion, we make use of trapezoidal velocity profiles for waypoint tracking.
-The maximum velocity for the trapezoid is set to the planned $\rho_{vg}$ and $\rho_{va}$ respectively, but the actual mean velocity is slightly less than the commanded due to the velocity profile.
-Finally, the maximum commanded speed has been limited to match the speed of the slowest subsystem, which in this implementation is the UGV.
-This has been done for safety reasons and to facilitate the coordination of the vehicles.
-In the experiments, we set the speeds of both the UGV and the UAV ($\rho_{vg}$ and $\rho_{va}$, respectively) to $0.25m/s$.
-4.4.2 UGV and UAV coordination
-In order to track the trajectory computed by our approach, the UGV and the UAV should follow the trajectories coordinately.
-This task is carried out by our Trajectory Coordination Module (TCM) (see Fig.
-7).
-In it, we opt for a loosely-coupled solution based on time synchronization.
+The maximum velocity for the trapezoid is set to the planned $\rho_{vg}$ and $\rho_{va}$, respectively, but the actual mean velocity is slightly less than the commanded due to the velocity profile.
+The maximum commanded speed has been limited to match the speed of the slowest subsystem, which in this implementation is the UGV.
+This decision was made for safety reasons and to facilitate the coordination of the vehicles.
+In the experiments, we set the speeds of both the UGV and the UAV ($\rho_{vg}$ and $\rho_{va}$, respectively,) to 0.25 m/s.
+Finally, in order to track the trajectory computed by our approach, the UGV and the UAV should follow the trajectories coordinately.
+This task is carried out by our TCM (see Figure 7).
+We opt for a loosely coupled solution based on time synchronization.
 TCM can only command new waypoints to both the UGV and UAV simultaneously, provided that both platforms have successfully reached the previous one.
 Thus, when a subsystem reaches its current waypoint, it waits for the other in case of need.
-5 Experiments
+5 Experimental Results
 We carried out different field experiments with our marsupial configuration in three different buildings of the UPO, Seville (Spain).
-Scenario 1 is located in Building 45, home of the Service Robotic Laboratory facilities.
+Scenario 1 is situated in Building 45, home to the Service Robotics Laboratory facilities.
 Next, we conducted additional tests in two old, abandoned buildings on the same university campus, both of which show significant structural damage.
-Scenario 2 is situated in the Abandoned Thermal Station, while Scenario 3 is located in the Old Theater building.
-Additional details of the experiments are provided in the video attached to this paper.
-5.1 Scenario 1: Flight Duration Tests
-The main goal of this experiment was to perform the necessary system integrity checks of the whole marsupial system and to estimate the flight endurance of the system in manual mode.
-To obtain a valid estimation, we mimicked the stop-and-go behavior in an inspection mission, obtaining a total trajectory length exceeding one kilometer on both the UAV and the UGV.
-Figure 12 shows two snapshots obtained during the experiment.
+Scenario 2 takes place within the Abandoned Thermal Station, whereas Scenario 3 is set in the Old Theater building.
+The logs of all experiments are accessible through our institutional repository (https://robotics.upo.es/datasets/marsupial, accessed on 3 Nov 2025).
+Furthermore, additional experimental details are presented in the video accompanying this paper.
+5.1 Scenario 1. Experiments 1–3: Flight Duration Tests
+The main goals of this experiment were to perform the necessary system integrity checks of the whole marsupial system and to estimate the flight endurance of the system.
+We performed three different experiments (Experiments 1–3) in which our marsupial system was commanded in manual mode, mimicking the stop-and-go behavior in an inspection mission.
+The duration of all the experiments exceeded one hour, with Experiment 3 exceeding two hours.
+All the experiments were performed in the same load condition.
+Figure 12 shows two snapshots obtained during the experiments.
 Refer to caption Refer to caption
-Figure 12: Two snapshots of the flight duration tests at building 45.
-The experiment exceeded one hour duration, which allowed us to get information about the discharging rate of both the onboard UAV and the batteries on the UGV used for powering the UAV via the Elistair tether system.
-Table 1 represents the level of the batteries powering the tether system and the onboard UAV battery used as a backup, for ease of comparison, during the experiment.
-In the second row, we can find the battery percentage of the onboard battery used for backup purposes on the UAV.
-Note that this battery slowly discharges until stabilizing at 86%, when the voltage level of the battery reaches the one provided by the Elistair system.
-Regarding the batteries onboard the UGV, their discharge rate remains constant during the experiment and very close together.
-At the end of the experiment, the remaining battery level was $53\%$, giving an average power consumption on the UAV of 1635 W.
-Table 1: Charge of the batteries, in [%] of the different systems as the mission time progresses.
-Battery Percentage
-Time [min] 0 15 30 45 60
-Onboard UAV Battery [%] 100 92 88 86 86
-Battery Pack 1 [%] 100 88 76 64 52
-Battery Pack 2 [%] 100 89 77 66 54
-5.2 Scenario 2: UAV Autonomous Navigation Test
-In Scenario 2, we visited the Abandoned Thermal Station building of the UPO.
-This building was built in the decade of the 1950s and presented great structural damage and debris on its surface that prevented our UGV ARCO platform from safely performing autonomous navigation on it.
-Therefore, we designed some experiments focused on the UAV system to test the indoor UAV localization and navigation systems.
-5.2.1 UAV Localization tests
-In a first set of experiments, we tested the ability of our localization system to provide us with an online estimation of the pose of the UAV in manual mode.
+Figure 12: Two snapshots of one flight duration test at building 45.
+Figure 13 represents the level of the batteries powering the LTS4 and the onboard UAV backup battery in Experiments 1–3.
+Remarkably, the discharge rate of the LTS4 battery remained approximately constant throughout all three experiments; the discharge is 46% in one hour in the worst case.
+Regarding the backup battery, the remaining level was above 80% in all three cases, which indicates that there was not significant usage, as the LTS4 system worked without issues during all three experiments.
+Moreover, this remaining battery level would allow the monitoring operator to safely land the UAV in the event of an LTS4 malfunction, as was designed in Section 2.4.
+With all the gathered data, we can conclude that our proposed system can perform inspection missions lasting up to two hours.
+Figure 13: Remaining battery levels of LTS4 and backup batteries in Experiments 1–3.
+5.2 Scenario 2: Abandoned Thermal Station
+Scenario 2 is located at the Abandoned Thermal Station building of the UPO, built in the decade of the 1950s and currently exhibiting significant structural damage and surface debris.
+These conditions prevented our ARCO platform from autonomously navigating within the building.
+Consequently, we designed a set of experiments focused on the UAV system to evaluate its indoor localization and navigation capabilities.
+5.2.1 Experiments 4–5: UAV Localization Tests
+In Experiments 4 and 5, we tested the precision of our localization system in manual mode.
 To this end, we designed a challenging experiment in which we made our UAV move around the whole experimental scenario at different height levels while recording the data from the sensors.
 Then, we tested different localization methods to compare their results and their average computation time.
-To evaluate the precision of the localization algorithm, we got the distance from each LiDAR detection inside a measurement to the closest obstacle in the map, to check for errors in the localization.
+To evaluate the precision of the localization algorithm, we obtained the distance from each LiDAR detection inside a measurement to the closest obstacle in the map to check for errors in the localization.
 The lower the average error in a measurement is, the better the vehicle is localized in the environment.
@@
 Table 3 represents the mean distance from the LiDAR points to the closest obstacle with different localization methods and their average execution time.
-Results indicate that our DLL method used in our system outperforms the MCL in terms of precision and runs significantly faster when compared to the ICP method, achieving similar errors.
-Moreover, the measured distance from the closest obstacle results indicate that the localization error should be in the order of centimeters, while also inferring a good precision in the estimation of the yaw.
-This precision is enough to safely perform cooperative missions in our marsupial system.
-Thus, DLL is a very convenient alternative for having precise localization when performing long-endurance inspection missions.
-Table 2: Average computation time of the update step on each method and mean distance from each point of the point-cloud (PC) to the closest obstacle in the map in Scenario 2.
-Method Avg. Computation time (s) Avg. PC distance to obstacle (m)
-AMCL-3D 0.54 0.45
-ICP 2.67 0.13
-DLL 0.08 0.14
-5.2.2 UAV autonomous flight tests
-Once we demonstrated that our localization system was able to obtain a precise estimation of the pose of the robot in real-time, we executed several autonomous tests for the UAV to test our trajectory tracking system.
-To this end, a simple inspection plan was designed and generated the trajectory to be tracked by the UAV by using our trajectory planner, assuming a static UGV.
-Figure 13 represents the generated trajectories to perform the UAV autonomous navigation test represented in the 3D model of the environment.
-Our system was able to autonomously follow the trajectory, obtaining mean lateral cross-track and height errors of 0.07 m and 0.08 m, respectively.
+We configured the MCL method with 500 particles, and we considered that each measure of the OS1 sensor has a standard deviation of 0.1 m and that the odometry has a relative error of 10%.
+We used the ICP implementation of the Point Cloud Library (PCL) with a maximum of 50 iterations, a maximum correspondence distance of 0.1 m and a RANSAC outlier rejection threshold of one meter.
+Finally, our DLL solver was implemented with the Ceres library, being configured to use 8 threads.
+All the experiments were executed in the NUC11TNKi5 computer onboard the UAV, which has a 13th Gen Intel® Core™ processor with 12 cores, one thread per core, and 16 GB of RAM.
+Results indicate that our DLL method outperforms the MCL in terms of precision and runs significantly faster when compared to the ICP method, achieving similar errors.
+While these results offer an indirect evaluation of the accuracy of the localization, in practice we found that it was enough to safely perform cooperative missions in our marsupial system, making DLL the better method for real-time localization.
+For more details about the experiment, please refer to the attached video, where you can check the behavior of our DLL in the localization tests.
+Table 3: Average computation time of the update step on each method and average distance from each point of the point-cloud to the closest obstacle in the map in Scenario 2.
+Exp. Method Comp. time (s) Avg. Dist. (m)
+MCL-3D 0.71 0.45
+4 ICP 1.8 0.13
+DLL 0.07 0.14
+MCL-3D 0.64 1.25
+5 ICP 2.9 0.56
+DLL 0.11 0.26
+5.2.2 Experiment 6: UAV Autonomous Flight Test
+After demonstrating that our localization system could accurately estimate the robot’s pose in real time, we conducted Experiment 6 to evaluate the UAV’s trajectory tracking system.
+To this end, a simple inspection plan was designed, and the trajectory to be tracked by the UAV was generated using our trajectory planner, assuming a static UGV.
+Figure 14 represents the generated trajectory to perform the UAV autonomous navigation test represented in the 3D model of the environment.
+The tracking system was able to autonomously follow the trajectory in all the executions, achieving mean lateral cross-track and altitude errors of 0.07 m and 0.08 m, respectively.
+Please refer to the attached video for additional details regarding the performance of the trajectory tracker.
 Figure 14: Reference trajectory (blue line) tracking experiments in Scenario 2.
-The position of the reel system is on a green dot and the red lines represent the tether configurations during the experiment.
-5.3 Scenario 3: Emulated Inspection Test
-Finally, to test the whole system in autonomous mode, we emulated a structural assessment inspection.
-In Scenario 3, our marsupial system is configured with a mission designed to look for emulated defects in the Old Theater building of the UPO.
-To emulate the defects, we disposed of an array of twelve AR markers with a side of fifteen centimeters on one of the walls of the building (see Fig.
-1b).
+The position of the reel system is on a green dot, and the red lines represent the tether configurations during the autonomous flight test.
+The points in white represent the 3D map of Scenario 2.
+5.3 Scenario 3. Experiment 7: Emulated Inspection Test
+Finally, to test the whole system in autonomous mode, we emulated a structural assessment inspection in Experiment 7, which was carried out in Scenario 3.
+In Experiment 7, our marsupial system is commanded to execute a mission designed to look for emulated defects in the Old Theater building of the UPO.
+To emulate the defects, we deployed an array of twelve Augmented Reality (AR) markers with a side of fifteen centimeters on one of the walls of the building (see Figure 1b).
 We equipped the UAV with a high-resolution camera to automatically detect the markers by using the ArUco library [43].
-Figure 8.a represents an obtained 3D map of the environment, which was used for localization and path planning purposes on both platforms in Scenario 3.
+Figure 8a represents an obtained 3D map of the environment, which was used for localization and path planning purposes on both platforms in Scenario 3.
 The goal of the experiment is to make an in-depth inspection of one of the walls of the building.
-To this end, we disposed of the PoIs of our inspection plan following a zig-zag line.
-Then, we generated a trajectory for the marsupial system using our trajectory planner module (see Section 4.3).
-In the experimental session, the batteries of the platform allowed us to perform the execution three times without the need of changing any batteries, for a total flight time of about fifty minutes.
-The trajectories commanded to the UAV and the UGV are represented in Fig.
-14.
-The UAV platform accurately tracked the trajectories, detecting all emulated defects in three experimental runs, as illustrated in Fig.
-15.
+To this end, we distributed the PoIs of our inspection plan following a zig-zag line.
+The distribution of the PoIs was sufficient to ensure complete wall coverage with a spatial resolution of twenty pixels per square centimeter, which is adequate for the detection of the emulated defects.
+We used the Fields2Cover library [44] to generate the sequence of PoIs (see Figure 15).
+Figure 15: Computed trajectory for the inspection mission of UGV (in green dots) and UAV (in blue dots).
+The square in white represents the initial UAV position.
+The spheres in white represent the PoIs to be visited.
+5.3.1 Trajectory Planning Results
+With the inspection plan defined as a sequence of PoIs, we generated a trajectory for the marsupial system using our trajectory planner module (see Section 4.3).
+Note that we need multiple calls to the planner to generate trajectories connecting the different PoIs.
+Table 4 presents the statistics of our trajectory planning method compared to the approach proposed in [28], which accounts for a non-straight tether when generating the complete inspection mission.
+Both methods successfully generated trajectories.
+Notably, our planning method achieves lower execution times compared to the baseline.
+Regarding clearance, our method yields a smaller mean distance between the tether and both the obstacles and the UAV.
+This is expected, as we consider only a straight tether, which reduces the available maneuvering space.
+Finally, in terms of total traveled distance, both methods produce similar UAV paths; however, our method requires the UGV to move slightly more due to the LoS constraint.
+In conclusion, our method is capable of generating safe trajectories for the system with lower computation time compared to the approach in [28].
+Table 4: Trajectory planning statistics for the proposed planner (LoS) and the planner proposed in [28] (catenary).
+Method Execution time (s) Avg. distance to obstacles (m) Trajectory Length (m)
+Cable UAV UGV UAV UGV
+Catenary 3.37 1.51 1.54 1.09 50.40 11.59
+LoS 0.49 1.11 1.08 1.65 48.30 14.14
+5.3.2 Mission Execution Results
+In the experimental session, our battery-powered LTS4 system allowed us to execute the inspection mission three times without the need of changing any batteries, for a total duration of thirty-five minutes.
+The trajectories commanded to the UAV and the UGV are represented in Figure 15 and are composed of 245 waypoints.
+As detailed in Section 2.4, a safety operator supervised the execution of the experiment, ensuring the absence of unexpected tether entanglements and monitoring the charge level of the UAV’s onboard backup battery.
+Figure 16 illustrates the battery levels of the UAV, the UGV, and the batteries supplying the LTS4 system.
+The backup battery level remained consistently above 80% throughout the experiment.
+Moreover, since no unexpected entanglements were observed, no operator intervention was required.
+Regarding the UGV’s onboard batteries, their voltage level remained approximately constant at 25.5 V during the entire experiment.
+Although voltage is not a precise indicator of the remaining charge for LiFePO4 batteries, this stability suggests that the battery operated within its mid-range (approximately 30 to 70%).
+Finally, the batteries powering the LTS4 maintained a charge level above 60%, indicating that safe operation could have continued for more than one hour after the experiment.
+The performance of our trajectory tracking systems was also remarkable.
+The system was capable of reaching all the planned waypoints in all three experimental runs.
+Finally, we evaluate the performance of the TCM.
+Table 5 presents the statistics of the waiting times for both the UAV and the UGV during each mission execution.
+The results are highly consistent.
+The UAV experiences a higher number of waiting events due to its lower velocity.
+However, the difference in waiting times is small, averaging approximately 1.5 s.
+Although this value may seem relatively high, it should be noted that the trajectory includes segments where only one vehicle is moving, causing the other to wait until the next waypoint is reached.
+Finally, the relatively high standard deviation indicates the presence of outliers in which the waiting times are prolonged for several seconds.
+This typically occurs at the beginning of the experiment.
+Figure 16: Remaining battery levels of the LTS4 (red and blue) and backup batteries and measured voltage of the UGV batteries (green and cyan) in Experiment 7.
+Table 5: Average waiting time and standard deviation for the UAV and UGV induced by the TCM on Experiment 7. Values in parentheses indicate the number of waiting events.
+Mission execution UAV waiting time (s) UGV waiting time (s)
+1 1.89 $\pm$ 2.21 (147) 1.51 $\pm$ 1.93 (147)
+2 1.43 $\pm$ 2.73 (144) 1.43 $\pm$ 2.14 (101)
+3 1.47 $\pm$ 2.39 (142) 1.39 $\pm$ 1.78 (103)
+The marsupial system was able to detect all the emulated defects in the three experimental runs, as illustrated in Figure 17.
 Statistics on the error in the estimated position of the detected ArUco markers are provided in Table 6.
@@
 Moreover, the cases with greater errors are usually related to a low number of observations of the AR marker.
-Figure 14: Computed trajectory for the inspection mission of UGV (in green dots) and UAV (in blue dots).
-The square and the sphere in white color represent the initial and the final positions, respectively.
-Figure 15: Current and past detections of emulated defects obtained in the inspection mission of Scenario 3.
+The achieved precision of the estimated position of the AR markers also validates the precision of our localization system.
+Figure 17: Current and past detections of emulated defects (AR Markers) obtained in Scenario 3, Experiment 7.
+(a) Current image from the onboard camera with an overlay of the current detection (ID 413).
+(b) Estimated positions of the UAV, the UGV and the current and past detections over the 3D map of the environment.
 Table 6: Average error and number of observations in the position estimation of the AR markers.
@@
 Observations 88 28 277 28 215 445 198 568 344 307 205 29
-6 Discussion and take-home messages
-This paper presents the design, implementation, and validation of a marsupial robotic system capable of performing long-duration inspections in GNSS-denied environments.
-Experimental results confirm the hypothesis that a tethered UAV-UGV system can significantly extend UAV flight time, opening the door to long-duration inspections.
+6 Discussion
+This paper has presented the design, implementation, and validation of a marsupial robotic system capable of performing long-duration inspections in GNSS-denied environments.
+Experimental results confirm the hypothesis that a tethered UAV-UGV system can extend UAV flight time to more than two hours of continuous operation, opening the door to long-duration inspections.
 The accuracy in detecting AR markers, with a mean error in the order of centimeters, demonstrates the viability of the system for detailed inspection tasks.
-Even though power consumption was higher than expected, the platform proved capable of performing inspection missions lasting more than two hours.
-The additional discharge may be due to extra connected components which were not included in the calculations, such as the LiDAR and the router, or due to heat losses in the different voltage conversions.
-However, the flight duration can be further extended by adding more battery packs to the UGV, as we are far away from its payload limits.
-A summary of the most valuable lessons we have learned from experimentation is mentioned in the Section 6.1.
-Then, Section 6.2 lists the advantages and disadvantages of our approach compared to other alternatives.
-6.1 Lessons Learned
-Several factors are essential for the success of a multi-robot experiment. In this section, we highlight some key insights that we have acquired from our experimentation.
-  • Design a setup procedure and a checklist to ensure experimental quality.
-When performing experiments with complex systems such as our tether marsupial system, it is important to carefully design a procedure for initializing and checking the systems before starting a mission.
-We follow the steps outlined in Algorithm 1.
-In Steps 1–3, the UGV is started system first, as its onboard computer hosts the ROS master node required for system-wide communication.
-Additionally, we power up the UAV as late as possible (Step 8) to maximize its available flight time.
-Conducting integrity tests on each onboard sensor is essential to avoid incomplete experiment recordings.
-Specifically, we verify that each sensor’s data rate and quality meet expected standards—this is done in Step 4 for the UGV and Step 10 for the UAV.
-  • Avoid dangerous configurations of the UAV, the UGV and the tether.
-The Elistair LIGH-T V4 station has a DC motor that performs a constant tension to the tether, keeping it taut.
-Experimentally, we found out that a tension of one Newton is enough to keep it straight in most situations.
-The main challenge here is that this tension introduces disturbances that must be managed by the UAV’s control system.
-When tension is applied downwards, it has a similar effect to increasing the system’s payload.
-However, when tension is applied laterally, the control system must adjust the UAV’s attitude to compensate for the disturbances.
-Therefore, it is advisable to penalize configurations with low elevation angles.
-Lastly, we secure the tether as close to the center of gravity of the UAV as possible to prevent the generation of torques.
-  • Prepare and label several UAV batteries the day before an experimental session.
-UAVs need a significant amount of energy to complete long-duration flights, which typically requires fresh batteries for each experimental run.
-However, LiPO batteries should not be stored fully charged for extended periods.
-Therefore, it is advisable to check and charge the UAV batteries the day before an experimental session.
-It also applies to our system because even if the energy consumption of the onboard UAV backup battery is low during the experiment (see Table 1), the M210 firmware expects the onboard batteries to be fully charged to power up the system.
-Algorithm 1 Experiment Setup and Execution Checklist
-1:Turn on the battery packs and the inverter at the UGV.
-2:Turn on router onboard the UGV.
-3:Turn on the UGV and start its ROS nodes, including the ROS master process necessary for communications.
-4:Turn on laptop at the base station and check for ROS connectivity with the UGV.
-5:Perform integrity tests of the sensors and software systems on the UGV.
-6:Set the UAV and the UGV at their initial poses.
-7:Turn on the Elistair Power Tether.
-8:Just before the experiment: Turn on the UAV computer.
-9:Check for connectivity to the UAV system
-10:Perform integrity tests to the sensors and software systems onboard the UAV.
-11:Start the TCM module on the base station.
-12:Start the recording of the experiments including video and data logging.
-13:Start the mission.
-6.2 Marsupial systems: benefits and drawbacks
-With all the knowledge gathered from experimentation, we are confident that our marsupial system is a viable alternative that has multiple advantages over existing approaches in the literature.
-Thus, the following list summarizes the benefits of our approach.
-  • The UGV can supply power to the UAV via the cable, significantly extending flight time compared to the UAV’s limited battery capacity. UAV time fly can extend over one hour.
-  • Unlike UAVs tethered to a fixed location, the UGV provides mobility, enabling the UAV to inspect a wider range of areas of interest.
-  • The UAV offers aerial mobility and monitoring capabilities, while the UGV provides a stable power source and additional payload capacity.
-  • The cable can serve as a reliable communication link and can also enhance safety by providing a physical connection between the two robots.
-  • Marsupial UAV-UGV systems have proven valuable in hazardous and confined environments, as exemplified by their use by the CSIRO team during the DARPA Subterranean Challenge.
-Our system has been able to carry out a long-endurance inspection task, in a GNSS-denied environment, with accurate localization and safe trajectory planning.
-However, even though our approach has been demonstrated as an effective alternative to perform long-duration inspection tasks, there are some disadvantages compared to other existing approaches.
-Besides, our approach has room for improvement in some areas.
-  • The tether can introduce significant disturbances in the UAV control subsystem. Interactions between the tether and the UAV generate dynamic forces that affect stability and control.
-  • The tether restricts the UAV’s range of motion and requires careful management to prevent collisions.
-Maintaining proper tension is crucial: downward tension impacts the system payload, while lateral tension requires UAV control adjustments to compensate for perturbations.
-  • There is a risk of the tether becoming entangled or snagged on obstacles.Therefore, the mission should be carefully planned to prevent the cable from touching any obstacles.
-In addition, the cable should also monitored during the experiments, which should be canceled in the event of entanglements or other unexpected events.
-  • Variable speed trajectory tracking can be implemented to smoothen the coordinated movement of the UAV and the UGV.
+The platform proved capable of performing inspection missions lasting more than two hours, which aligns with our estimates when accounting for the power consumption of the UGV’s internal LiDAR and router, both of which are connected to the inverter.
+However, the flight duration could be further extended by integrating more battery packs to the UGV, as the platform remains well below its payload limit.
+With the knowledge gathered from our experimentation, we are confident that the proposed marsupial system represents a viable alternative with several advantages over existing approaches in the literature.
+One of its key strengths lies in the capability of the UGV to supply power to the UAV through the tether, which significantly extends the UAV’s endurance beyond the limitations imposed by its onboard battery.
+In practice, this enables flight times of over two hours, a remarkable improvement over conventional untethered UAVs.
+Furthermore, in contrast to systems where the UAV is tethered to a fixed station, our mobile UGV base provides additional flexibility by enabling the UAV to cover a broader range of inspection areas.
+The complementary nature of both platforms enhances mission capabilities: while the UAV contributes aerial mobility and monitoring functionality, the UGV acts as a stable power source and offers additional payload capacity.
+Marsupial UAV–UGV systems have already demonstrated their utility in hazardous and confined environments, as shown, for instance, by the CSIRO team in the DARPA Subterranean Challenge [19] and the Oxpecter platform [21].
+Our approach significantly advances autonomy, demonstrating long-endurance inspection in GNSS-denied environments with precise localization and safe trajectory planning.
+Despite these benefits, our approach is not without limitations.
+While it is true that tethering limits the UAV’s maneuverability and may constrain its flyable space, requiring careful management to prevent collisions, the extended flight duration enables our solution to carry out long-duration inspection missions, where steady flight is essential to ensure the validity of the measurements.
+In our view, the extended flight time outweighs the reduction in maneuverability and speed for the applications considered.
+Maintaining proper tension is essential, since downward tension directly affects payload capacity, whereas lateral forces demand corrective control actions from the UAV to maintain stability.
+However, we found that by configuring LTS4 to perform a tension of 1 N to the UAV, we can keep the cable straight when only increasing the required force to hover in a 2.
+5%, as the UAV weighs approximately four kilograms.
+Another drawback is that the tether carries the risk of becoming entangled or snagged on environmental obstacles.
+To mitigate this, missions are carefully designed to prevent any contact between the tether and the environment.
+The cable is continuously monitored during the experiments, which are halted if entanglements or other unexpected events occur.
+Finally, while our system achieves effective cooperation between the UAV and UGV, further improvements could be obtained through variable-speed trajectory tracking, which would enable smoother coordination between both platforms.
 7 Conclusions
-This paper presents our solution for long-duration inspection: a tethered marsupial system connecting a UAV and a UGV.
-In it, we detail the hardware and software configurations that allow autonomous navigation in UAV inspection tasks.
-Notably, the system achieved over one hour of flight time—and up to a maximum of two hours—thanks to the power supplied by the batteries onboard the UGV.
+This paper presents our solution for long-duration inspection: a tethered marsupial system connecting a UAV to a UGV.
+In it, we detail the hardware and software configurations that allow autonomous navigation in UAV inspection tasks, as demonstrated in several field experiments.
+Notably, we have demonstrated that the proposed system has a flight duration of over two hours thanks to the power supplied by the batteries onboard the UGV.
 Therefore, the proposed marsupial system successfully extends the operational capabilities of the UAV.
-The experimental results obtained across three different scenarios validate the robustness and operational capabilities of our system in real-world conditions, delivering reliable inspection performance in challenging, GNSS-denied environments.
-Specifically, a series of tests conducted at the University Pablo de Olavide confirmed the system’s ability to carry out autonomous, long-duration inspection missions.
-Notably, our implementation of the DLL localization system provided real-time pose estimations with centimeter-level accuracy, outperforming state-of-the-art methods.
-Furthermore, our path tracking system successfully followed the reference trajectories, maintaining maximum tracking errors below ten centimeters.
-As a result, the system was able to detect simulated defects with centimeter-level accuracy during the emulated inspection test.
+The proposed tethered marsupial system can successfully perform autonomous inspection tasks, detecting targets such as markers without being hindered by the tether.
+The inspection tasks are carried out with high accuracy both in terms of the localization of the marsupial robot and in the detection of the markers.
+Future Work
 Future work will focus on refining the power system to further extend the operational range.
-Additionally, we are developing a continuous control system within the TCM module to enable smooth coordination between platforms by adjusting the tracking velocity of each system, moving away from the current Stop & Go approach.
-Moreover, we are interested in expanding the inspection capabilities of the system by integrating additional sensors required for in-depth inspection tasks, such as radar and IR cameras.
+Additionally, we are developing a continuous control system within the TCM to enable smooth coordination between platforms by adjusting the tracking velocity of each system, moving away from the current stop-and-go synchronization.
+Moreover, we aim to expand the inspection capabilities of the system by integrating additional sensors required for advanced inspection tasks, such as radar and IR cameras.
 However, the UAV’s payload limitations may prevent it from carrying all sensors simultaneously.
-To address this, we are taking two key actions.
+To address this, we are pursuing two complementary approaches.
 First, we are designing modular sensor attachments that allow easy reconfiguration of the UAV based on the specific requirements of each inspection mission.
-Second, Elistair offers a tethered station which includes an optical fiber communication link.
-In this way, we could remove the PC onboard the UAV, making room for installing additional sensors.
-Declarations
-[…]
-Funding This work was supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21, funded by MCIN/AEI/ 10.13039/501100011033 and ”European Union NextGenerationEU/PRTR”.
-[…]
-Conflict of interest/Competing interests The authors declare that they have no conflicts of interest or competing interests.
-[…]
-Ethics Approval Not Applicable.
-[…]
-Consent for publication Not Applicable.
-[…]
-Consent to participate Not Applicable.
-[…]
-Data availability
-Data will be available upon reasonable request. Also, the code is available on the GitHub organization "robotics-upo" [32].
-[…]
-Authors’ Contributions Simón Martínez-Rozas and David Alejo: Conceptualization, data curation, software, investigation, methodology, analysis, experiments, writing original draft.
-José Javier Carpio: Mechanical development and maintenance of the robot, experiments, writing-review and editing.
-Fernando Caballero and Luis Merino: Conceptualization, validation, resources, supervision, funding acquisition, project management, writing-review and editing.
-[…]
+Second, Elistair offers a tethered station that includes an optical fiber communication link.
+In this way, we could remove the computer onboard the UAV, making room for additional sensors.
+Another area of interest is the investigation of the applicability of our system in outdoor scenarios, or even in mixed scenarios that encompass both indoor and outdoor areas.
+These mixed scenarios pose significant challenges for the localization system, as they involve partial GNSS coverage and require ensuring a smooth transition between GNSS and GNSS-denied localization modes.
+In addition, the environmental factors such as wind can adversely affect the performance of the control system.
+Finally, we aim to also use this approach in emergency response situations.
+To this end, we plan to modify the traction type of our UGV platform to enable safe and reliable navigation over uneven terrain.
+Additionally, we plan to include a multi-robot SLAM method to perform autonomous exploration of unknown scenarios.
+All the data captured in the experiments is available at https://robotics.upo.es/datasets/marsupial, accessed on 3 Nov 2025.
 Acknowledgements
-We would like to thank the staff of the Service Robotics Laboratory at the Universidad Pablo de Olavide for their unselfish help during the experiments.
+We would like to thank the staff of the Service Robotics Laboratory at UPO for their unselfish help during the experiments.
 We would also like to give special thanks to Sara Victoria Hernández Díaz for suggesting different experimental scenarios for structural inspection and granting us access to them.
+This work was partially supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21, funded by MCIN/AEI/ 10.13039/501100011033 and “European Union NextGenerationEU/PRTR”.